### v1

# Highly Efficient Self-Adaptive Reward Shaping for Reinforcement Learning

Abstract
Reward shaping addresses the challenge of sparse rewards in reinforcement learning by constructing denser and more informative reward signals. To achieve self-adaptive and highly efficient reward shaping, we propose a novel method that incorporates success rates derived from historical experiences into shaped rewards. Our approach utilizes success rates sampled from Beta distributions, which dynamically evolve from uncertain to reliable values as more data is collected. Initially, the self-adaptive success rates exhibit more randomness to encourage exploration. Over time, they become more certain to enhance exploitation, thus achieving a better balance between exploration and exploitation. We employ Kernel Density Estimation (KDE) combined with Random Fourier Features (RFF) to derive the Beta distributions, resulting in a computationally efficient implementation in high-dimensional continuous state spaces. This method provides a non-parametric and learning-free approach. The proposed method is evaluated on a wide range of continuous control tasks with sparse and delayed rewards, demonstrating significant improvements in sample efficiency and convergence stability compared to several baselines.
1 Introduction
Environments with sparse and delayed rewards present notable challenges for reinforcement learning (RL), despite its significant progress in various fields. In such contexts, as the reward structure lacks immediate signals, agents receive feedback only after long horizons, making the ability to quickly and effectively discover beneficial samples crucial for successful learning (Ladosz et al., 2022). To address this, a straightforward solution is to reconstruct the reward models by introducing auxiliary signals to assess the agent’s behavior at each step, which has led to the popular technique of Reward Shaping (RS) (Sorg et al., 2010a, b; Gupta et al., 2022). Inverse reinforcement learning (IRL), which extracts reward functions from human knowledge or expert demonstrations, represents an intuitive approach within this framework. However, this heavily relies on extensive human input and data, which can be difficult to obtain especially in complex environments. Alternatively, fully autonomous approaches have emerged as an attractive direction.
Automatically maintained shaped rewards can be generally categorized into two branches: intrinsic motivation based signals, which are task-agnostic, and hidden value based signals, which are commonly task-related. The former mainly introduces exploration bonuses to encourage agents to explore a wider range of states, typically by rewarding novel or less frequently visited states (Ostrovski et al., 2017; Devidze et al., 2022; Tang et al., 2017; Burda et al., 2018; Bellemare et al., 2016). These methods have shown effective exploration enhancement, however, often overlooking the internal values of the states themselves. This can lead to the "noisy TV" problem, where agents become fixated on highly novel but meaningless regions, thus trapping agents in suboptimal behaviors (Mavor-Parker et al., 2022). In contrast, the latter leverages high-level heuristics to guide agents in retrieving hidden values from their experiences, which can significantly improve the agent’s convergence (Trott et al., 2019; Ma et al., 2024; Zheng et al., 2018).
To overcome the limitations of existing RS methods, and try to combine the advantages of exploration-encouraged and hidden value-based rewards, this paper proposes a Self-Adaptive Success Rate based reward shaping mechanism (SASR) [^1] The source code is accessible at: https://github.com/mahaozhe/SASR. The success rate, defined as the ratio of a state’s presence in successful trajectories to its total occurrences, works as an auxiliary reward distilled from historical experience. The success rate assesses a state’s contribution toward successful task completion, which aligns with the agent’s original objectives, offering informative guidance for agent learning. Furthermore, to mitigate overconfidence caused by deterministic success rates, we adopt Beta distributions to model success rates from a probabilistic perspective. Beta distributions allow for a self-adaptive evolution of confidence in approximating a state’s success rate. This effectively prevents premature certainty and ensures stable convergence to reliable rewards as more data accumulates. To derive Beta distributions, we leverage kernel density estimation enhanced by random Fourier features to efficiently estimate success and failure counts, formulating a highly efficient approach. The main contributions of this paper are summarized as follows:
  1. (i)
    We propose SASR, an autonomous reward-shaping mechanism without the requirement of expert demonstrations or domain knowledge. By extracting a metric from historical experiences that aligns with the agent’s optimization objectives, SASR effectively densifies the environmental rewards.
  2. (ii)
    We introduce a novel self-adaptive strategy to achieve a balance between exploration and exploitation. Initially, low-confidence Beta distributions provide stochastic rewards that encourage exploration by diversifying the policy optimization directions. With the accumulation of experience, high-confidence Beta distributions offer more reliable and precise rewards to enhance exploitation.
  3. (iii)
    To derive the Beta distributions in continuous state spaces, we implement kernel density estimation, leveraging random Fourier features. This formulates a non-parametric and learning-free approach, thereby achieving remarkably low computational complexity.
  4. (iv)
    We use sparse-reward and continuous control environments, MuJoCo (Todorov et al., 2012) and simulated robot (de Lazcano et al., 2023), to evaluate the performance of SASR. It significantly outperforms several baselines in terms of sample efficiency, learning speed, and convergence stability.
2 Related Work
Reward shaping (RS) techniques can be broadly categorized based on the source of learning: either from human knowledge or agent’s own experiences. Techniques drawing reward models from human knowledge rely on pre-existing expertise, where methods such as Inverse Reinforcement Learning (IRL) (Arora and Doshi, 2021; Ramachandran and Amir, 2007; Ziebart et al., 2008; Hadfield-Menell et al., 2016) and Inverse Optimal Control (IOC) (Schultheis et al., 2021) aim to derive reward or objective functions from expert demonstrations. Another branch within this approach, Imitation Learning (IL) (Ho and Ermon, 2016; Duan et al., 2017; Brown et al., 2020), attempts to clone human behaviors from offline data without explicitly modeling reward structures. Following this, transferring the extracted reward models to new tasks has received considerable efforts (Bıyık et al., 2022; Wu et al., 2021; Ellis et al., 2021; Cheng et al., 2021; Adamczyk et al., 2023). However, these methods require significant human-driven data and generally show limited adaptability to out-of-distribution (OOD) scenarios. Therefore, we focus instead on the autonomous self-learning approaches, which based on the meaning of rewards, can be discussed from two perspectives: intrinsic motivation based and hidden value based rewards.
Intrinsic motivation based RS explores general heuristics or task-agnostic metrics to encourage exploration in sparse-reward environments. Potential-based algorithms define the shaped reward in the form of $\gamma\Phi(s^{\prime})-\Phi(s)$, where $\Phi(\cdot)$ is a potential function, enabling the reward to be nullified in the Bellman equation, thus ensuring the optimal policy remains invariant (Devlin and Kudenko, 2012; Asmuth et al., 2008). However, the potential function’s design heavily depends on environmental dynamics, making it more suitable in model-based RL. More popular methods incorporate exploration bonuses by rewarding novel states (Devidze et al., 2022; Hong et al., 2018; Eysenbach et al., 2019). For instance, count-based strategies maintain visitation counts and assign higher rewards to less frequently visited states. To overcome the challenge of state counting in continuous spaces, Tang et al. (2017) utilized a hash function to discretize the state space, Bellemare et al. (2016) introduced pseudo-counts based on recording probabilities, and Ostrovski et al. (2017) used PixelCNN Van den Oord et al. (2016). Our work adopts a KDE approach to estimate the state counts. Additionally, Burda et al. (2018) proposed random network distillation to assess state novelty in high-dimensional spaces; Curiosity-driven approaches reward encounters with states that are surprising and out-of-prediction (Sun et al., 2022; Pathak et al., 2017). The exploration bonuses branch shows effective exploration enhancement, but only considering novelty while ignoring the underlying values may lead to sub-optimal behaviors. A notable example illustrating this limitation is the "noisy TV" problem (Mavor-Parker et al., 2022), where agents may become fixated on irrelevant and dynamic stimuli that yield high novelty but no useful information.
Hidden value based RS, on the other hand, primarily focuses on task-related signals to reveal the extent to which states contribute to achieving higher environmental rewards and their underlying significance. Trott et al. (2019) integrated the distance between a state and the target as additional rewards. Stadie et al. (2020) derived informative reward structures using a Self-Tuning Network (STN) to optimize reward guidance. Memarian et al. (2021) captured the preferences among different trajectories by ranking them via a trained classifier. Zheng et al. (2018) minimized the KL-divergence between the learned and the original environmental rewards to synchronize their distributions. Mguni et al. (2023) used an auxiliary agent that competed against the original agent in a Markov game; Instead, Ma et al. (2024) introduced ReLara, a collaborative framework where an assistant reward agent automatically generates rewards to guide the policy agent. Moreover, the incorporation of multiple agents or hierarchical structures to share and transfer knowledge through synchronized reward functions is also a popular research direction (Park et al., 2023; Yi et al., 2022; Gupta et al., 2023; Hu et al., 2020).
While intrinsic motivation rewards directly encourage exploration, hidden value rewards support efficient convergence, Ma et al. (2024) combined these two advances by introducing a scheme that evolves the shaped rewards from random perturbations to meaningful metrics. Injecting random rewards during early training relaxes the assumption that only novel states should be rewarded, suggesting instead that any state can be randomly rewarded. Building on this concept, our study considers the success rate, which is the probability that a state will contribute to task completion, as the hidden value. We sample this metric from evolving Beta distributions that progressively increase their confidence, thus aligning the process to balance the exploration and exploitation adaptively. Compared with ReLara, which relies on a black-box agent, our method derives the shaped rewards directly from historical data, effectively reducing the delay of reward convergence.
3 Preliminaries
Reinforcement Learning (RL) aims to train an agent to interact with an environment, which is commonly modeled as a Markov Decision Process (MDP). An MDP represented as $\langle S,A,T,R,\gamma\rangle$, involves four main components: $S$ is the state space, $A$ is the action space, $T:S\times A\times S\rightarrow[0,1]$ is the probability of transitioning from one state to another given a specific action, and $R:S\rightarrow\mathbb{R}$ is the reward model. The objective in RL is to learn a policy $\pi(a|s)$ that maximizes the expected cumulative rewards $G=\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t})]$, where $\pi(a|s)$ indicates the likelihood of selecting action $a$ in state $s$, and $\gamma$ is the discount factor. Model-free RL methods include updating value functions, directly optimizing policies, or employing an actor-critic framework (Sutton and Barto, 2018).
Beta Distribution is defined on the interval $[0,1]$, making it ideal for modeling proportions or probabilities. It is parameterized by $\alpha$ and $\beta$, which are interpreted as prior counts of successes and failures of a binary outcome. The probability density function of a Beta-distributed variable $X$ is:
$f(x;\alpha,\beta)=\frac{1}{B(\alpha,\beta)}x^{\alpha-1}(1-x)^{\beta-1},$ (1)
where $B(\alpha,\beta)$ is a normalization factor. The key characteristic of Beta distribution is its adaptability: as more data accumulates, the values of $\alpha$ and $\beta$ increase, the distribution’s shape becomes narrower, thereby increasing confidence in the estimated probabilities. This feature is particularly beneficial in adaptive online learning, aligning closely with our objective of balancing exploration and exploitation.
Kernel Density Estimation (KDE) is a non-parametric approach to approximate the probability density function of a random variable from data samples. Given $n$ data points $\{x_{i}\}_{i=1}^{n}$, KDE smooths these points to approximate the density function as follows:
$\hat{d}(x)=\frac{1}{nh}\sum_{i=1}^{n}K\left(\frac{x-x_{i}}{h}\right),$ (2)
where $h$ is the bandwidth, $K(\cdot)$ is a kernel function such as Gaussian kernel, Laplacian kernel, or Cauchy kernel. KDE is particularly useful in scenarios where the actual distribution is not well-defined or complex, such as continuous state spaces in our RL environments.
4 Methodology
We propose a Self-Adaptive Success Rate based reward shaping mechanism (SASR) to accelerate RL algorithms in sparse-reward environments. Figure 1 illustrates the principles of the SASR mechanism: The diagram consists of two parts representing the early and late stages of learning. From the early stage to the late stage, the Beta distributions modeling the estimation of the true success rates evolve from being more stochastic to more deterministic as experiences accumulate. This progression closely aligns with the agent’s exploration-exploitation balance. The process of Beta distributions’ evolution and generation of shaped rewards from them are detailed in Section 4.1. Additionally, to achieve highly efficient computation, we leverage KDE and RFF to estimate success and failure counts, which are used to derive the corresponding Beta distributions, as detailed in Section 4.2. Lastly, Section 4.3 presents the integration of SASR into the RL agent and the overall algorithmic flow.
Figure 1: A schematic diagram of the self-adaptive success rate based reward shaping mechanism. KDE: Kernel Density Estimation; RFF: Random Fourier Features.
4.1 Self-Adaptive Success Rate Sampling
We consider the following augmented reward function for SASR:
$R^{SASR}(s)=R^{E}(s)+\lambda R^{S}(s),$ (3)
where $R^{E}(s)$ is the environmental reward, $R^{S}(s)$ is the additional shaped reward, and $\lambda$ is a weight factor to control its relative scale. When $\lambda\rightarrow 0$, we can recover the original reward model. We assign the shaped reward $R^{S}(s_{i})$ of a given state based on its success rate – defined as the ratio of the state’s presence in successful trajectories to its total occurrences. This metric provides a meaningful and informative reward from a statistical perspective, as a higher success rate indicates a greater likelihood and contribution for the state to guide the agent towards successful task completion, thus deserving a higher shaped reward. Formally, the success rate based shaped reward $R^{S}(s_{i})$ is given by:
$R^{S}(s_{i})=f\left(\frac{N_{S}(s_{i})}{N_{S}(s_{i})+N_{F}(s_{i})}\right),$ (4)
where $N_{S}(s_{i})$ and $N_{F}(s_{i})$ denote the counts of appearances for state $s_{i}$ in historical successful and failed trajectories respectively, and $f(\cdot)$ is a mapping function that adjusts the magnitude of rewards: $f:[0,1]\rightarrow[R_{\min},R_{\max}]$.
Given the counts $N_{S}(s_{i})$ and $N_{F}(s_{i})$ derived from prior experiences, directly using a deterministic success rate $\frac{N_{S}(s_{i})}{N_{S}(s_{i})+N_{F}(s_{i})}$ may lead to overconfidence in the estimation of the true value. To address this, we model the success rate estimation from a probabilistic perspective, inspired by the principles of Thompson sampling, an effective algorithm widely used to balance exploration and exploitation in multi-armed bandit problems (Thompson, 1933; Agrawal and Goyal, 2012). We model the success rate of each state as an approximated variable in a Beta distribution, setting the shape parameters $\alpha=N_{S}(s_{i})+1$ and $\beta=N_{F}(s_{i})+1$:
$r^{S}_{i}\sim\text{Beta}(r;\alpha,\beta)=\frac{1}{B(N_{S}(s_{i})+1,N_{F}(s_{i})+1)}r^{N_{S}(s_{i})}(1-r)^{N_{F}(s_{i})},$ (5)
where the beta function $B(\cdot,\cdot)$ is the normalization factor. By sampling from this distribution, we obtain a probabilistic estimate of the state’s success rate. This sampled value, $r^{S}_{i}$, is then processed through the mapping function $f(\cdot)$ to produce the shaped reward: $R^{S}(s_{i})=f(r^{S}_{i})$.
As $N_{S}(s_{i})$ and $N_{F}(s_{i})$ progressively increase throughout the learning process, they influence the shape and sampling variability of the Beta distribution. Generating the shaped reward from these evolving Beta distributions offers several advantages:
  • Encourage Exploration. In the early stages of learning, the counts $N_{S}(s_{i})$ and $N_{F}(s_{i})$ are relatively low, resulting in broader Beta distributions and greater variability in sampled values, which injects significant randomness into the shaped rewards. These random rewards optimize the agent in various directions with tiny steps, thus shifting the anchors from which the stochastic policy samples actions. The strategy effectively broadens the range of policies and actions, thus encouraging the exploration. (Supporting experimental results are provided in Section 5.2.)
  • Enhance Exploitation. In the late stage of learning, as the counts $N_{S}(s_{i})$ and $N_{F}(s_{i})$ increase, the Beta distribution gradually sharpens, concentrating generated rewards closer to the true success rate. The evolution to a certain reward signal with higher confidence highly supports the agent’s exploitation phase, facilitating more efficient convergence towards optimal policies.
  • Consistent Optimization. The peak of the Beta distribution, $B(N_{S}(s_{i})+1,N_{F}(s_{i})+1)$, is given by $\frac{N_{S}(s_{i})}{N_{S}(s_{i})+N_{F}(s_{i})}$, directly approximating the success rate. Meanwhile, the expected value of the distribution, $\frac{N_{S}(s_{i})+1}{N_{S}(s_{i})+N_{F}(s_{i})+2}$, aligns closely with the calculated success rate. This ensures that, despite some randomness, the overall reward remains consistent with policy optimization.
4.2 Highly Efficient Beta Distribution Derivation
In this section, we present the methods to obtain the success and failure counts, $N_{S}(s_{i})$ and $N_{F}(s_{i})$, which are required for deriving the Beta distributions. In tabular environments with low-dimensional, discrete, and finite state spaces, tracking these counts can be straightforwardly accomplished using two tables. However, most real-world environments involve high-dimensional, continuous, and infinite state spaces, making it impractical to maintain exact counts for every state.
To address this challenge, we use Kernel Density Estimation (KDE) to approximate the densities of successes and failures from accumulated experience. Specifically, we maintain two buffers, $\mathcal{D}_{S}$ and $\mathcal{D}_{F}$, to store the states in successful and failed trajectories, respectively. By treating these states as scattered data instances distributed across the state space, KDE estimates the density as:
$\tilde{d}_{X}(s_{i})=\frac{1}{|\mathcal{D}_{X}|}\sum_{j=1}^{|\mathcal{D}_{X}|}{K(s_{i}-s_{j})},\quad X\in\{S,F\},$ (6)
where $K(\cdot)$ is the kernel function, and $|\mathcal{D}_{X}|$ is the total number of states in the corresponding buffer. We select the Gaussian kernel in our implementation. The estimated density $\tilde{d}_{X}(s_{i})$ serves as a proxy for the likelihood of encountering state $s_{i}$ under success or failure scenarios, providing a statistically sound basis for estimating $N_{X}(s_{i})$. By multiplying $\tilde{d}_{X}(s_{i})$ with the total number of observed states $N$, and representing states as $k$-dimensional vectors, the count $\tilde{N}_{X}(s_{i})$ is estimated as:
$\tilde{N}_{X}(\bm{s}_{i})=N\times\tilde{d}_{X}(\bm{s}_{i})=\frac{N}{|\mathcal{D}_{X}|}\sum_{j=1}^{|\mathcal{D}_{X}|}{\exp\left(-\frac{||\bm{s}_{i}-\bm{s}_{j}||^{2}}{2\sigma^{2}}\right)},\quad X\in\{S,F\},$ (7)
where $\sigma$ is a hyperparameter that controls the bandwidth of the Gaussian kernel.
We further integrate Random Fourier Features (RFF) (Rahimi and Recht, 2007) to reduce computational complexity, as calculating the Gaussian kernel can be expensive, especially in scenarios involving high-dimensional state spaces and large buffers. RFF approximates the kernel function of the original $k$-dimensional states through an inner product of $M$-dimensional randomized features:
$K(\bm{s}_{i},\bm{s}_{j})\approx z(\bm{s}_{i})^{T}z(\bm{s}_{j}),\quad z(\bm{s})=\sqrt{\frac{2}{M}}\cos(\bm{W}^{T}\bm{s}+\bm{b}),$ (8)
where $z(\cdot)$ is the RFF mapping function with $\bm{W}=\big{[}\bm{w}^{(1)},\ldots,\bm{w}^{(M)}\big{]}\in\mathbb{R}^{k\times M}$ and $\bm{b}=\big{[}b^{(1)},\ldots,b^{(M)}\big{]}^{T}\in\mathbb{R}^{M}$ randomly sampled from the following spectral distributions:
$\bm{w}^{(m)}\sim\mathcal{N}(\bm{0},\sigma^{-2}\bm{I}_{k}),\quad b^{(m)}\sim\text{Uniform}(0,2\pi),\quad m=1,\ldots,M,$ (9)
where $\bm{I}_{k}$ is the $k\times k$ identity matrix. Equation 9 is applied for the Gaussian kernel, while different kernels, such as Laplacian or Cauchy kernels, have different spectral distributions. Detailed derivations of the RFF method and the list of spectral distributions are provided in Appendix A.
Suppose for each iteration, the buffer size of $\mathcal{D}_{X}$ is $D$ and the batch size is $B$, the computational complexity to retrieve the corresponding count $N_{X}$ is $O(MDB)$, representing a linear complexity (detailed derivation in Appendix B). RFF transforms the nonlinear kernel computations into vectors’ linear operations, significantly accelerating computation by leveraging the vectorization capabilities of GPUs (Dongarra et al., 2014). This demonstrates its higher efficiency compared to other methods that rely on network updates and inferences involving extensive nonlinear computations.
Moreover, we introduce a hyperparameter, the retention rate $\phi\in(0,1]$, to regulate the volume and diversity of states stored in the buffers. Instead of storing all the states in each trajectory, we uniformly retain a specific portion $\phi$. This strategy offers two main benefits: (1) as adjacent states in one trajectory tend to be highly similar, especially those closer to the initial state are repetitive and uninformative, retaining a portion of states can skip redundant states and increase sample diversity; (2) maintaining a lower retention rate in the early training stage keeps lower $N_{S}$ and $N_{F}$, resulting in a broader Beta distribution and avoiding premature overconfidence.
Given the retention rate $\phi$, the space complexity of maintaining two buffers is $O(\phi T|s|)$, where $T$ is the total number of iterations and $|s|$ is the size of a single state. To address scalability concerns in more complex or high-dimensional scenarios, we propose leveraging the existing replay buffer of the backbone algorithm, soft actor-critic. By augmenting this buffer with a flag array that indicates whether each state is a success or failure, while this introduces additional indexing computations during processing, we can efficiently save space requirements.
4.3 The SASR Mechanism for RL agents
Algorithm 1 Self-Adaptive Success Rate based Reward Shaping
0:  Environment $\mathcal{E}$ and agent $\mathcal{A}$.
0:  Experience replay buffer $\mathcal{D}$.
0:  State buffers for success $\mathcal{D}_{S}$ and failure $\mathcal{D}_{F}$.
0:  RFF mapping function $z:\mathbb{R}^{k}\rightarrow\mathbb{R}^{M}$. $\triangleright$ Sample $\bm{W}$ and $\bm{b}$ based on Equation 9
1:  for each trajectory $\tau=\emptyset$ do
2:     for each environmental step do
3:        $(s_{t},a_{t},s_{t+1},r^{E}_{t})\leftarrow\text{CollectTransition}(\mathcal{E},\mathcal{A})$triangleright interact with the environment
4:        $\mathcal{D}\leftarrow\mathcal{D}\cup\{(s_{t},a_{t},s_{t+1},r^{E}_{t})\}$triangleright store the transition in the replay buffer
5:        $\tau\leftarrow\tau\cup\{s_{t}\}$triangleright record the state in the trajectory
6:     end for
7:     if trajectory is successful: $\mathcal{D}_{S}\leftarrow\mathcal{D}_{S}\cup\tau$triangleright store the trajectory in the success buffer
8:     else: $\mathcal{D}_{F}\leftarrow\mathcal{D}_{F}\cup\tau$triangleright otherwise, store the trajectory in the failure buffer
9:  end for
10:  for each update step do
11:     $\{(s_{t},a_{t},r^{E}_{t},s_{t+1})_{i}\}\sim\mathcal{D}$triangleright sample a batch of transitions from the replay buffer
12:     $\tilde{N}_{S}=z(\bm{s}_{t})^{T}z(\bm{\mathcal{D}_{S}})$triangleright estimate the success counts
13:     $\tilde{N}_{F}=z(\bm{s}_{t})^{T}z(\bm{\mathcal{D}_{F}})$triangleright estimate the failure counts
14:     $r^{S}_{t}\sim\text{Beta}(r;\tilde{N}_{S}+1,\tilde{N}_{F}+1)$triangleright sample the success rate from the Beta distribution
15:     $r^{SASR}_{t}=r^{E}_{t}+\lambda f(r^{S}_{t})$triangleright compute the SASR reward
16:     Update agent $\mathcal{A}$ with $\{(s_{t},a_{t},r^{SASR}_{t},s_{t+1})_{i}\}$
17:  end for
Building upon the SASR reward, we employ the soft actor-critic (SAC) algorithm of Haarnoja et al. (2018a) as the foundation for our RL agent. Let $\pi^{\mathcal{A}_{P}}_{\theta}$ be parameterized policy network and $Q^{\mathcal{A}_{P}}_{\phi}$ be parameterized Q-network. We optimize the Q-function by minimizing the mean squared error (MSE) between the predicted Q-value and the temporal difference (TD) target:
$\mathcal{L}(\phi)=\Big{(}Q_{\phi}(s_{t},a_{t})-\big{(}r^{E}_{t}+\lambda R^{S}(s_{t})+\gamma Q_{{\phi}^{\prime}}(s_{t+1},a_{t+1})\big{)}\Big{)}^{2},$ (10)
where $Q_{{\phi}^{\prime}}$ is obtained from a secondary frozen target network to maintain a fixed objective (Mnih et al., 2015). It is worth noting that the environmental reward $r^{E}_{t}$ is retrieved from the replay buffer. Conversely, the shaped reward $R^{S}(s_{t})$ is computed in real-time using the most recently updated $N_{S}(s_{t})$ and $N_{F}(s_{t})$, thereby capturing the latest advancements in learning progress.
We optimize the policy network by maximizing the expected Q-value and the entropy of the policy $\mathcal{H}\big{(}\pi^{\mathcal{A}_{P}}(\cdot|s_{t})\big{)}$ (Haarnoja et al., 2018b):
$\mathcal{L}(\theta)=\mathbb{E}_{a_{t}\sim\pi_{\theta}(\cdot|s_{t})}\big{[}-Q_{\phi}(s_{t},a_{t})+\log{\pi_{\theta}(a_{t}|s_{t})}\big{]}.$ (11)
The outline of the SASR is summarized in Algorithm 1. We maintain two buffers to store successful and failed states and estimate the success and failure counts from them. Subsequently, shaped rewards are computed by sampling from Beta distributions followed by a scale transformation. The shaped rewards, combined with environmental rewards, are used to update the agent’s Q-function and policy.
In summary, SASR offers two primary advantages: First, it leverages the adaptive properties of Beta distributions to continuously adjust the agent’s confidence levels in understanding the environment, effectively maintaining the balance between exploration and exploitation. Second, SASR operates as a non-parametric, learning-free approach, enabling highly efficient reward calculation.
5 Experiments
We evaluate SASR on continuous control environments, including seven MuJoCo tasks (Todorov et al., 2012): AntStand, AntSpeed, AntFar, AntVerFar, WalkerKeep, HumanStand, and HumanKeep, two robotic tasks (de Lazcano et al., 2023): RobotReach and RobotPush, as well as one physical simulation task MountainCar (Towers et al., 2023), shown in Figure 2. Our environment setup follows the guidelines established in the ReLara paper (Ma et al., 2024), with ReLara also included as a baseline for comparison. The tasks provide only sparse rewards, with a reward of $1$ granted upon reaching the final objective within the maximum permitted steps, and 00 for all other states. To ensure robust experimental validation, we train multiple instances for each setting using different random seeds and report the average results. Additionally, we maintain uniform hyperparameters and network architectures across all experiments, as detailed in Appendix D.
AntStand
AntSpeed
AntFar
AntVeryFar
WalkerKeep
HumanKeep
HumanStand
RobotReach
RobotPush
MountainCar
Figure 2: MuJoCo, robotic and physical simulation tasks with their objectives in our experiments.
5.1 Comparison and Discussion
Baselines. We compare SASR with eight baselines to benchmark its performance: (a) RL with an Assistant Reward Agent (ReLara) (Ma et al., 2024), (b) RL Optimizing Shaping Algorithm (ROSA) (Mguni et al., 2023), (c) Exploration-Guided RS (ExploRS) (Devidze et al., 2022), (d) Count-based static hashing exploration (#Explo) (Tang et al., 2017), (e) Random Network Distillation (RND) (Burda et al., 2018), (f) Soft Actor-Critic (SAC) (Haarnoja et al., 2018a), (g) Twin Delayed Deep Deterministic Policy Gradient (TD3) (Fujimoto et al., 2018), and (h) Proximal Policy Optimization (PPO) (Schulman et al., 2017). Algorithms (a) to (e) are all reward shaping methods, involving either exploration bonuses or auxiliary agents to provide additional rewards to tackle the sparse-reward challenge, while algorithms (f) to (h) are advanced RL algorithms. For the implementation of these baselines, we employ either the CleanRL library (Huang et al., 2022) or codes provided by the authors.
Figure 3 and Table 1 report the average episodic returns and standard errors throughout training. Our findings indicate that SASR surpasses the baselines in terms of sample efficiency, learning stability, and convergence speed. The primary challenges in these environments are attributed to delayed and sparse rewards alongside long horizons, making exploration essential for timely obtaining successful trajectories. Although the exploration strategies of algorithms such as ExploRS, #Explo, and RND are designed to reward novel states, effectively expanding the early exploration space with the direct additional target, they continue focusing on discovering novel states, overlooking the implicit values of these states, which makes them fail to return to the final objectives.
Figure 3: The learning performance of SASR compared with the baselines.
Table 1: The average episodic returns and standard errors of SASR and the baselines.
Algorithms AntStand AntSpeed AntFar AntVeryFar WalkerKeep HumanStand HumanKeep RobotReach RobotPush MountainCar
SASR 39.12$\pm$2.86 0.94$\pm$0.07 73.92$\pm$4.97 78.64$\pm$3.92 158.24$\pm$5.59 42.63$\pm$2.17 180.98$\pm$4.40 81.29$\pm$6.52 137.06$\pm$12.66 0.91$\pm$0.04
ReLara 28.66$\pm$1.82 0.33$\pm$0.02 67.77$\pm$4.30 64.07$\pm$4.17 77.14$\pm$8.77 29.72$\pm$1.85 160.31$\pm$7.30 103.56$\pm$7.18 58.71$\pm$6.98 0.89$\pm$0.01
ROSA 3.80$\pm$0.03 0.02$\pm$0.00 4.71$\pm$0.28 0.64$\pm$0.07 32.14$\pm$1.19 8.55$\pm$0.03 152.38$\pm$4.98 0.27$\pm$0.03 $0$0.90$\pm$0.02
ExploRS 4.52$\pm$0.04 0.01$\pm$0.00 5.42$\pm$0.22 1.67$\pm$0.10 2.47$\pm$0.13 8.63$\pm$0.03 158.09$\pm$4.42 0.79$\pm$0.04 0.20$\pm$0.08 -0.99$\pm$0.02
#Explo 6.79$\pm$0.50 0.00$\pm$0.00 12.40$\pm$1.66 0.55$\pm$0.09 131.56$\pm$5.40 28.73$\pm$1.79 160.60$\pm$7.04 4.19$\pm$0.42 6.31$\pm$0.85 0.79$\pm$0.02
RND 4.23$\pm$0.03 0.02$\pm$0.00 6.66$\pm$0.25 1.00$\pm$0.09 44.78$\pm$1.39 8.67$\pm$0.03 159.79$\pm$4.27 28.18$\pm$2.53 0.04$\pm$0.04 0.94$\pm$0.00
SAC 15.93$\pm$0.69 0.00$\pm$0.00 9.64$\pm$0.57 20.73$\pm$2.87 70.96$\pm$8.10 9.31$\pm$0.05 4.59$\pm$0.84 45.03$\pm$4.92 0.55$\pm$0.21 -0.05$\pm$0.02
TD3 0.00$\pm$0.00 0.00$\pm$0.00 0.81$\pm$0.02 0.81$\pm$0.02 18.62$\pm$0.75 5.72$\pm$0.04 0.55$\pm$0.03 0.00$\pm$0.00 0.00$\pm$0.00 0.00$\pm$0.00
PPO 4.54$\pm$0.04 0.00$\pm$0.00 6.33$\pm$0.24 1.80$\pm$0.11 33.77$\pm$1.11 8.36$\pm$0.03 138.13$\pm$12.64 79.52$\pm$10.80 0.00$\pm$0.00 0.93$\pm$0.00
SASR, on the other hand, utilizes the success rate as auxiliary information to assess each state’s contribution to task completion. This metric is consistent with the agent’s primary objectives, and incorporating this meaningful signal into the temporal difference updating at each step significantly enhances learning efficiency. Without this integration, the agent would have to wait until the end of the trajectory to obtain informative signals. Furthermore, SASR regulates an efficient exploration strategy by injecting substantial random rewards during the early phases to optimize the agent in multiple directions thereby increasing the likelihood of collecting positive samples. This is achieved by sampling from initially wider Beta distributions. Importantly, these Beta distributions adaptively yield increasingly confident and accurate success rates as more data is collected, thereby accelerating the agent’s learning. These strategies collectively enhance SASR’s efficiency and stability in handling these challenging tasks.
While ReLara employed a similar exploration mechanism with SASR by rewards perturbations to encourage exploration, it relies on a separate black-box agent, which requires more iterations to converge. In contrast, SASR’s success rate sampling is more intuitive and direct, reducing the delay in providing valuable information to the policy agent. ReLara’s advantage lies in its integration of the policy agent’s action into the reward construction, as evidenced in the RobotReach task. The target point for the robot is randomly selected in each episode, this can be considered as a higher dimension for SASR, because of the changing target. In this case, ReLara achieves better results, with its access to action information. This can be improved in SASR by incorporating the action as an additional feature in the state vector.
5.2 Effect of Self-Adaptive Success Rate Sampling
SASR introduces a novel self-adaptive mechanism that balances exploration and exploitation by maintaining the randomness of the shaped rewards. Figure 4 illustrates the generated rewards at five stages of learning in the AntStand task. Figure 4 (bottom) represents the learning curve, and Figure 4 (top) illustrates the distribution of the generated rewards over the "height of the ant" feature, one of the dimensions in the state vector.
Figure 4: Distribution of the shaped rewards over the height of the ant robot in the AntStand task, at different stages of the training process.
As learning progresses, the evolving shaped rewards exhibit two crucial attributes: the values transition from random to meaningful, and the variance decreases from uncertain to deterministic. In the early stages of training, the shaped rewards involve significant randomness due to higher uncertainty. Although these random signals contain limited information, they prompt the agent to take tiny optimization steps in diverse directions, effectively shifting the anchors of the policies to various regions. This strategy broadens the actions generated by SAC’s stochastic policy, encouraging the agent’s exploration and increasing the diversity of collected samples. From another perspective, the perturbations on the rewards can be seen as a relaxation of the assumption that only novel states should be rewarded, instead, randomly reward any state, which also shows its empirical effectiveness. In the later stages, the rewards are observed to converge to closely match the height of the robot, a meaningful and intuition-aligned metric, which enhances the agent’s exploitation of its policy. This self-adaptive mechanism for balancing exploration and exploitation improves SASR’s ability to gather positive samples earlier, thereby facilitating a faster convergence to the optimal policy.
5.3 Ablation Study
We conduct ablation studies to investigate five key components of SASR: (1) sampling from Beta distributions versus using the deterministic success rates; (2) the retention rate $\phi$, which controls the volume of historical information; (3) the bandwidth $h$ of the kernel density estimation; (4) the number of dimensions $M$ of RFF features; and (5) the weight factor $\lambda$, which adjusts the scales of the shaped rewards. We select six representative tasks to report the experimental results, and the quantitative data is provided in Appendix C.
Sampling from Beta distributions. We examine a variant of SASR that omits sampling from the Beta distribution and instead directly uses the success rate $N_{S}(s_{i})/(N_{S}(s_{i})+N_{F}(s_{i}))$. Comparative results are shown in Figure 5. In the early stages of learning, limited experience makes this success rate an inaccurate estimate of the true value. Using this overly confident fixed value can mislead the agent. Additionally, removing the process of sampling from the Beta distribution disrupts the agent’s exploration encouraged by random rewards, leading to slower convergence. The results highlight the importance of sampling from the Beta distribution for effective learning.
Figure 5: The impact of omitting Beta distribution sampling on the performance of SASR.
Figure 6: Learning performance of SASR on different retention rates $\phi$.
Retention rate $\phi$. The retention rate controls the counts’ estimation, subsequently impacting the Beta distribution and sampled rewards. We assess various retention rates in Figure 6. A high retention rate, such as $\phi=1$, preserving all samples, results in a densely populated and redundant experience pool, causing the Beta distribution to be prematurely overconfident, which can degrade performance. Conversely, a lower retention rate $\phi=0.01$ requires more iterations to accumulate sufficient samples, thereby slowing the convergence to meaningful rewards and affecting the agent’s convergence speed. The results suggest that carefully balancing the retention rate is crucial, and promising future work could involve adaptive adjustment of the retention rate during the learning process.
Bandwidth $h$ of Gaussian kernel. We evaluate different bandwidths $h$ of the Gaussian kernel for density estimation, as shown in Figure 7. Beyond fixed bandwidths, we also test a linearly decreasing bandwidth configuration ($h:0.5\rightarrow 0.1$), which represents progressively increasing confidence in KDE. Results indicate that an optimal bandwidth selection is crucial. A small bandwidth $h=0.01$ results in farther distances between samples, leading many to have zero estimated density; while a large bandwidth $h=1$ causes the samples indistinguishable due to the overly flat kernel function. Both scenarios lead to suboptimal performance. The decreasing bandwidth setting does not yield significant improvement and tends to reduce stability due to inconsistent density estimations.
Figure 7: Learning performance of SASR on different bandwidths $h$ of Gaussian kernel.
Figure 8: Learning performance of SASR on different RFF feature dimensions $M$.
Figure 9: Learning performance of SASR on different shaped reward weight factors $\lambda$.
RFF feature dimensions $M$. Learning performance on different RFF feature dimensions $M$ is shown in Figure 8. SASR shows relatively low sensitivity on $M$, provided it is large enough to capture the complexity of the states. Results indicate that values like $M=500,1000,2000$ all yield comparable performance levels. However, a significantly lower dimension, such as $M=50$, degrades performance, which is also suggested in the original RFF study (Rahimi and Recht, 2007).
Shaped reward weight factor $\lambda$. We evaluate the impact of different shaped reward scales on learning performance, as shown in Figure 9. We observe that SASR performs better with intermediate values like $\lambda=0.4,0.6,0.8$. At $\lambda=0.2$, the minimal shaped reward scale reduces state differentiation, leading to suboptimal performance. At $\lambda=1$, aligning the shaped reward scale with the environmental reward introduces excessive uncertainty and randomness, which may overwhelm the feedback disturbances and hinder learning. The findings highlight that maintaining a balanced reward scale is important for achieving optimal learning outcomes.
6 Conclusion and Discussion
We propose SASR in this paper, a self-adaptive reward shaping algorithm based on success rates, to address the sparse-reward challenge. SASR achieves an exploration and exploitation trade-off mechanism by generating shaped rewards from evolving Beta distributions. Experiments demonstrate that this adaptability significantly enhances the agent’s convergence speed. Additionally, the implementation of KDE and RFF formulates a highly efficient, non-parametric, and learning-free approach to deriving Beta distributions. This mechanism also provides a sound alternative to traditional count-based RS strategies, adapting effectively to continuous environments. Our evaluations confirm the superior performance of SASR in terms of sample efficiency and learning stability, and its robustness across various continuous-control tasks.
While SASR is designed for environments with sparse rewards, in settings where the environmental rewards are dense, the introduction and maintenance of additional shaped rewards could be unnecessary. On the other hand, SASR operates on the premise that sparse rewards indicate task completion to align the success rate notion. This correlation is generally straightforward, yet it may not intuitively apply in all contexts. Exploring how to extend SASR to more general sparse and even dense reward scenarios presents a promising direction for further research. Moreover, the derivation of Beta distributions heavily relies on the samples in the success and failure buffers. Currently, our method does not capture the relationships and varying importance of different states within the same trajectory, making the algorithm relatively sensitive to the retention rate, a crucial hyperparameter for achieving promising performance. Therefore, developing an adaptive retention rate or better mechanisms for maintaining the two buffers are crucial avenues for future enhancements.
Appendix A Derivation of Random Fourier Features
We incorporate Random Fourier Features (RFF) [Rahimi and Recht, 2007] to approximate the kernel functions in the KDE process for the SASR algorithm. Let the original state be $k$-dimensional, denoted as $\bm{s}\in\mathbb{R}^{k}$, and the kernel function be $k(\bm{s}_{i},\bm{s}_{j})$. RFF approximates the kernel function by projecting the input $k$-dimensional space into a $M$-dimensional feature space using a mapping function $z:\mathbb{R}^{k}\rightarrow\mathbb{R}^{M}$. The RFF-based kernel function is then defined as follows:
$k(\bm{s}_{i},\bm{s}_{j})\approx z(\bm{s}_{i})^{T}z(\bm{s}_{j}),$ (12)
We provide the derivation of the RFF approximation in this section.
First, we clarify that RFF primarily targets shift-invariant kernels, that satisfy $k(\bm{s}_{i},\bm{s}_{j})=k(\bm{s}_{i}-\bm{s}_{j})$. Common shift-invariant kernels include Gaussian kernels, Laplacian kernels, and Cauchy kernels. Given a shift-invariant kernel function $k(\Delta)$, we perform the inverse Fourier transform:
[…]
$=\mathbb{E}_{\bm{w}}\big{[}e^{\mathsf{i}\bm{w}^{T}(\bm{s}_{i}-\bm{s}_{j})}\big{]},$ (14)
where we can consider $\bm{w}\sim p(\bm{w})$ based on the Bochner’s theorem, and $p(\bm{w})$ is called the spectral distribution of kernel function. For the three types of shift-invariant kernels, the corresponding spectral distributions are listed in Table 2:
Table 2: Some shift-invariant kernels and their associated spectral distributions.
Kernel Kernel function, $k(\bm{s}_{i}-\bm{s}_{j})$ Spectral density, $p(\bm{w})$
Gaussian $\exp\left(-\frac{\|\bm{s}_{i}-\bm{s}_{j}\|_{2}^{2}}{2}\right)$ $(2\pi)^{-\frac{M}{2}}\exp\left(-\frac{\|\bm{w}\|_{2}^{2}}{2}\right)$
Laplacian $\exp\big{(}-\|\bm{s}_{i}-\bm{s}_{j}\|_{1}\big{)}$ $\prod_{m=1}^{M}\frac{1}{\pi(1+w_{d}^{2})}$
Cauchy $\prod_{i=1}^{k}\frac{2}{\pi\big{(}1+(\bm{s}_{i}-\bm{s}_{j})^{2}\big{)}}$ $\exp\big{(}-\|\bm{w}\|_{1}\big{)}$
Next, we perform the Euler’s formula transformation, which retains only the cosine term since we are dealing with real-valued functions, the kernel function can be further derived as:
$k(\bm{s}_{i},\bm{s}_{j})$ $=\mathbb{E}_{\bm{w}}\big{[}e^{\mathsf{i}\bm{w}^{T}(\bm{s}_{i}-\bm{s}_{j})}\big{]}$ (15)
$=\mathbb{E}_{\bm{w}}\big{[}\cos(\bm{w}^{T}(\bm{s}_{i}-\bm{s}_{j}))\big{]}$ (16)
$=\mathbb{E}_{\bm{w}}\big{[}\cos(\bm{w}^{T}(\bm{s}_{i}-\bm{s}_{j}))\big{]}+\mathbb{E}_{\bm{w}}\big{[}\mathbb{E}_{b}[\cos(\bm{w}^{T}(\bm{s}_{i}+\bm{s}_{j})+2b)]\big{]}$ (17)
$=\mathbb{E}_{\bm{w}}\big{[}\mathbb{E}_{b}[\cos(\bm{w}^{T}(\bm{s}_{i}-\bm{s}_{j}))+\cos(\bm{w}^{T}(\bm{s}_{i}+\bm{s}_{j})+2b)]\big{]}$ (18)
$=\mathbb{E}_{\bm{w}}\big{[}\mathbb{E}_{b}[\sqrt{2}\cos(\bm{w}^{T}\bm{s}_{i}+b)\sqrt{2}\cos(\bm{w}^{T}\bm{s}_{j}+b)]\big{]},$ (19)
where $b\sim\text{Uniform}(0,2\pi)$. Equation 17 holds since $\mathbb{E}_{b\sim\text{Uniform}(0,2\pi)}\big{[}\cos(t+2b)\big{]}=0$ for any $t$. Equation 19 is obtained from $\cos(A-B)+\cos(A+B)=2\cos(A)\cos(B)$, where $A=\bm{w}^{T}\bm{s}_{i}+b$, $B=\bm{w}^{T}\bm{s}_{j}+b$.
We define the mapping $z_{\bm{w},b}(\bm{s})=\sqrt{2}\cos(\bm{w}^{T}\bm{s}+b)$, then the kernel function can be approximated by the inner product of two vectors and the expectation can be approximated by Monte Carlo sampling:
$k(\bm{s}_{i},\bm{s}_{j})$ $=\mathbb{E}_{\bm{w}}\big{[}\mathbb{E}_{b}[z_{\bm{w},b}(\bm{s}_{i})z_{\bm{w},b}(\bm{s}_{j})]\big{]}$ (20)
$\approx\frac{1}{M}\sum_{m=1}^{M}z_{\bm{w}_{d},b_{d}}(\bm{s}_{i})z_{\bm{w}_{d},b_{d}}(\bm{s}_{j})$ (21)
$=z(\bm{s}_{i})^{T}z(\bm{s}_{j}).$ (22)
Therefore, we have derived the mapping function $z(\bm{s})=\sqrt{2/M}\cos(\bm{W}^{T}\bm{s}+\bm{b})$, where $\bm{W}\in\mathbb{R}^{M\times k}$ and $\bm{b}\in\mathbb{R}^{M}$. The RFF-based kernel function can be approximated by the inner product of the mapped features in the $M$-dimensional space.
Appendix B Derivation of Computation Complexity
In this section, we derive the computational complexity to retrieve the success or failure counts $N_{S}$ and $N_{F}$ for each iteration. Suppose the buffer size of $\mathcal{D}_{X}$ is $D$, the batch size of $\mathcal{B}$ is $B$, the corresponding counts are retrieved by calculating:
[…]
where the mapping function is defined as:
$z(\bm{s})=\sqrt{\frac{2}{M}}\cos(\bm{W}^{T}\bm{s}+\bm{b}),\quad\bm{W}\in\mathbb{R}^{k\times M},\quad\bm{b}\in\mathbb{R}^{M}.$ (24)
For each state, the mapping function calculation involves:
  • Matrix multiplication $\bm{W}^{T}\bm{s}$: $kM$.
  • Addition $\bm{W}^{T}\bm{s}+\bm{b}$: $M$.
  • Cosine calculation $\cos(\bm{W}^{T}\bm{s}+\bm{b})$: $M$.
Therefore, the computational complexity for calculating $z(\bm{s})$ for one state is $O(kM)$.
For each pair of states $(\bm{s}_{i},\bm{s}_{j})$, calculating the kernel involves $M$ multiplications and $M-1$ additions, thus, the complexity is $O(M)$.
For each iteration, we calculate the RFF mapping for all states in the buffer and the batch and then compute the kernel between them. The complexities involve three parts: RFF mapping for the buffer, RFF mapping for the batch and kernel calculation:
$O(DkM)+O(BkM)+O(MDB).$ (25)
Since the first two terms $O(DkM)$ and $O(BkM)$ are dominated by the last term $O(MDB)$ when the buffer size and the batch size are large, the overall computational complexity to retrieve the corresponding counts can be approximated as $O(MDB)$.
Appendix C Supplementary Experimental Results for Ablation Study
We provide the detailed quantitative results of the ablation study in this section.
Table 3: Ablation study #1: The average episodic returns and standard errors of SASR and the variant without sampling from Beta distributions.
Tasks SASR (with sampling) SASR (without sampling)
AntStand $\bm{39.12\pm 2.86}$ $23.25\pm 2.12$
AntFar $\bm{73.92\pm 4.97}$ $37.86\pm 5.16$
HumanStand $\bm{42.63\pm 2.17}$ $9.30\pm 0.083$
HumanKeep $\bm{180.98\pm 4.40}$ $170.51\pm 10.68$
RobotReach $\bm{81.29\pm 6.52}$ $33.76\pm 4.41$
RobotPush $\bm{137.06\pm 12.66}$ $70.37\pm 11.97$
Table 4: Ablation study #2: The average episodic returns and standard errors of SASR with different retention rates.
Tasks $\phi=1$ $\phi=0.1$ (default) $\phi=0.01$
AntStand $17.98\pm 1.86$ $\bm{39.12\pm 2.86}$ $20.27\pm 2.08$
AntFar $19.95\pm 3.00$ $\bm{73.92\pm 4.97}$ $48.47\pm 5.28$
HumanStand $9.25\pm 0.08$ $\bm{42.63\pm 2.17}$ $19.17\pm 1.52$
HumanKeep $179.10\pm 7.06$ $\bm{180.98\pm 4.40}$ $180.13\pm 6.27$
RobotReach $54.07\pm 8.39$ $\bm{81.29\pm 6.52}$ $35.18\pm 5.95$
RobotPush $3.91\pm 1.22$ $\bm{137.06\pm 12.66}$ $1.88\pm 0.47$
Table 5: Ablation study #3: The average episodic returns and standard errors of SASR with different bandwidths $h$ of Gaussian kernel.
Tasks $h=0.01$ $h=0.1$ $h=0.2$ (default) $h=1$ $h=0.5\rightarrow 0.1$
AntStand $7.94\pm 0.28$ $16.77\pm 1.18$ $\bm{39.12\pm 2.86}$ $12.76\pm 0.40$ $19.00\pm 1.63$
AntFar $11.68\pm 0.96$ $44.31\pm 3.34$ $\bm{73.92\pm 4.97}$ $27.96\pm 2.46$ $72.71\pm 4.57$
HumanStand $9.31\pm 0.06$ $\bm{43.64\pm 1.94}$ $42.63\pm 2.17$ $9.32\pm 0.06$ $13.81\pm 0.93$
HumanKeep $178.51\pm 4.99$ $\bm{181.44\pm 4.73}$ $180.98\pm 4.40$ $178.22\pm 4.97$ $178.04\pm 5.12$
RobotReach $37.03\pm 4.19$ $72.30\pm 5.67$ $\bm{81.29\pm 6.52}$ $35.87\pm 4.57$ $49.47\pm 5.1$
RobotPush $0.40\pm 0.13$ $8.80\pm 4.44$ $\bm{137.06\pm 12.66}$ $0.12\pm 0.06$ $0.34\pm 0.13$
Table 6: Ablation study #4: The average episodic returns and standard errors of SASR with different RFF feature dimensions $M$.
Tasks $M=50$ $M=500$ $M=1000$ (default) $M=2000$
AntStand $3.56\pm 0.12$ $12.08\pm 1.36$ $39.12\pm 2.86$ $\bm{42.52\pm 3.14}$
AntFar $27.76\pm 3.24$ $36.31\pm 2.69$ $\bm{73.92\pm 4.97}$ $61.79\pm 4.63$
HumanStand $9.46\pm 0.05$ $25.42\pm 1.83$ $42.63\pm 2.17$ $\bm{52.70\pm 2.13}$
HumanKeep $178.51\pm 4.84$ $181.81\pm 4.60$ $\bm{180.98\pm 4.40}$ $180.58\pm 4.63$
RobotReach $65.39\pm 6.22$ $38.25\pm 4.37$ $\bm{81.29\pm 6.52}$ $45.98\pm 4.90$
RobotPush $33.58\pm 4.86$ $54.56\pm 6.66$ $\bm{137.06\pm 12.66}$ $45.15\pm 6.15$
Table 7: Ablation study #5: The average episodic returns and standard errors of SASR with different shaped reward weight factors.
Tasks $\lambda=0.2$ $\lambda=0.4$ $\lambda=0.6$ (default) $\lambda=0.8$ $\lambda=1.0$
AntStand $7.60\pm 0.73$ $17.25\pm 1.58$ $\bm{39.12\pm 2.86}$ $28.43\pm 2.07$ $2.82\pm 0.09$
AntFar $22.51\pm 2.60$ $34.04\pm 3.87$ $\bm{73.92\pm 4.97}$ $67.62\pm 4.89$ $18.05\pm 1.79$
HumanStand $9.53\pm 0.04$ $33.95\pm 1.90$ $\bm{42.63\pm 2.17}$ $30.65\pm 1.65$ $12.23\pm 0.42$
HumanKeep $\bm{182.57\pm 4.32}$ $180.22\pm 4.60$ $180.98\pm 4.40$ $179.74\pm 4.53$ $179.62\pm 4.70$
RobotReach $41.05\pm 4.17$ $\bm{82.33\pm 6.37}$ $81.29\pm 6.52$ $71.30\pm 6.54$ $15.89\pm 2.06$
RobotPush $9.18\pm 1.99$ $118.92\pm 10.79$ $\bm{137.06\pm 12.66}$ $65.51\pm 12.71$ $47.98\pm 9.64$
Appendix D Network Structures and Hyperparameters
D.1 Network Structures
Figure 10 illustrates the structures of the policy network and Q-network employed in our experiments. The agent utilizes simple multilayer perceptron (MLP) models for these networks. Given the use of stochastic policies, the policy network features separate heads to generate the means and standard deviations of the inferred normal distributions, which are then used to sample actions accordingly.
Figure 10: The structures of policy network and Q-network in our implementation.
D.2 Hyperparameters
We have observed that SASR demonstrated high robustness and was not sensitive to hyperparameter choices. Table 8 shows the set of hyperparameters that we used in all of our experiments.
Table 8: The hyperparameters used in the SASR algorithm.
Hyperparameters Values
reward weight $\lambda$ (default) 0.6
kernel function bandwidth 0.2
random Fourier features dimension $M$ 1000
retention rate $\phi$ (default) 0.1
discounted factor $\gamma$ 0.99
replay buffer size $|\mathcal{D}|$ $1\times 10^{6}$
batch size 256
actor module learning rate $3\times 10^{-4}$
critic module learning rate $1\times 10^{-3}$
SAC entropy term factor $\alpha$ learning rate $1\times 10^{-4}$
policy networks update frequency (steps) $2$
target networks update frequency (steps) $1$
target networks soft update weight $\tau$ $5\times 10^{-3}$
burn-in steps $5000$
Appendix E Compute Resources
The experiments in this paper were conducted on a computing cluster, with the detailed hardware configurations listed in Table 9. The computing time for the SASR algorithm in each task (running 1,000,000 steps) was approximately $6\pm 2$ hours.
Table 9: The compute resources used in the experiments
Component Specification
Operating System (OS) Ubuntu 20.04
Central Processing Unit (CPU) 2x Intel Xeon Gold 6326
Random Access Memory (RAM) 256GB
Graphics Processing Unit (GPU) 1x NVIDIA A100 20GB
Brand Supermicro 2022

### v3

@@
 Abstract
-Reward shaping addresses the challenge of sparse rewards in reinforcement learning by constructing denser and more informative reward signals.
-To achieve self-adaptive and highly efficient reward shaping, we propose a novel method that incorporates success rates derived from historical experiences into shaped rewards.
-Our approach utilizes success rates sampled from Beta distributions, which dynamically evolve from uncertain to reliable values as more data is collected.
-Initially, the self-adaptive success rates exhibit more randomness to encourage exploration.
-Over time, they become more certain to enhance exploitation, thus achieving a better balance between exploration and exploitation.
-We employ Kernel Density Estimation (KDE) combined with Random Fourier Features (RFF) to derive the Beta distributions, resulting in a computationally efficient implementation in high-dimensional continuous state spaces.
-This method provides a non-parametric and learning-free approach.
-The proposed method is evaluated on a wide range of continuous control tasks with sparse and delayed rewards, demonstrating significant improvements in sample efficiency and convergence stability compared to several baselines.
+Reward shaping is a technique in reinforcement learning that addresses the sparse-reward problem by providing more frequent and informative rewards.
+We introduce a self-adaptive and highly efficient reward shaping mechanism that incorporates success rates derived from historical experiences as shaped rewards.
+The success rates are sampled from Beta distributions, which dynamically evolve from uncertain to reliable values as data accumulates.
+Initially, the shaped rewards exhibit more randomness to encourage exploration, while over time, the increasing certainty enhances exploitation, naturally balancing exploration and exploitation.
+Our approach employs Kernel Density Estimation (KDE) combined with Random Fourier Features (RFF) to derive the Beta distributions, providing a computationally efficient, non-parametric, and learning-free solution for high-dimensional continuous state spaces.
+Our method is validated on various tasks with extremely sparse rewards, demonstrating notable improvements in sample efficiency and convergence stability over relevant baselines.
 1 Introduction
-Environments with sparse and delayed rewards present notable challenges for reinforcement learning (RL), despite its significant progress in various fields.
-In such contexts, as the reward structure lacks immediate signals, agents receive feedback only after long horizons, making the ability to quickly and effectively discover beneficial samples crucial for successful learning (Ladosz et al., 2022).
-To address this, a straightforward solution is to reconstruct the reward models by introducing auxiliary signals to assess the agent’s behavior at each step, which has led to the popular technique of Reward Shaping (RS) (Sorg et al., 2010a, b; Gupta et al., 2022).
-Inverse reinforcement learning (IRL), which extracts reward functions from human knowledge or expert demonstrations, represents an intuitive approach within this framework.
-However, this heavily relies on extensive human input and data, which can be difficult to obtain especially in complex environments.
+Environments with extremely sparse rewards present notable challenges for reinforcement learning (RL).
+In such contexts, as the reward model lacks immediate signals, agents receive feedback only after long horizons, making the ability to quickly discover beneficial samples crucial for successful learning (Ladosz et al.,, 2022).
+To address this, a straightforward solution is to reconstruct the reward models by introducing auxiliary signals to assess the agent’s behavior, which has led to the popular technique of Reward Shaping (RS) (Strehl and Littman,, 2008; Gupta et al.,, 2022).
+Inverse reinforcement learning (IRL), which extracts reward functions from human knowledge or expert demonstrations, represents an intuitive approach within this framework Arora and Doshi, (2021).
+However, IRL heavily relies on extensive human input, which can be difficult to obtain especially in complex environments.
 Alternatively, fully autonomous approaches have emerged as an attractive direction.
-Automatically maintained shaped rewards can be generally categorized into two branches: intrinsic motivation based signals, which are task-agnostic, and hidden value based signals, which are commonly task-related.
-The former mainly introduces exploration bonuses to encourage agents to explore a wider range of states, typically by rewarding novel or less frequently visited states (Ostrovski et al., 2017; Devidze et al., 2022; Tang et al., 2017; Burda et al., 2018; Bellemare et al., 2016).
-These methods have shown effective exploration enhancement, however, often overlooking the internal values of the states themselves.
-This can lead to the "noisy TV" problem, where agents become fixated on highly novel but meaningless regions, thus trapping agents in suboptimal behaviors (Mavor-Parker et al., 2022).
-In contrast, the latter leverages high-level heuristics to guide agents in retrieving hidden values from their experiences, which can significantly improve the agent’s convergence (Trott et al., 2019; Ma et al., 2024; Zheng et al., 2018).
-To overcome the limitations of existing RS methods, and try to combine the advantages of exploration-encouraged and hidden value-based rewards, this paper proposes a Self-Adaptive Success Rate based reward shaping mechanism (SASR) [^1] The source code is accessible at: https://github.com/mahaozhe/SASR.
+Automatically maintained reward shaping can be broadly categorized into two branches: intrinsic motivation-based rewards, which are task-agnostic, and inherent value-based rewards, which are typically task-specific.
+The former mainly introduces exploration bonuses to encourage agents to explore a wider range of states, commonly by rewarding novel or infrequently visited states (Burda et al.,, 2018; Ostrovski et al.,, 2017; Tang et al.,, 2017; Bellemare et al.,, 2016).
+While these methods effectively enhance exploration, they tend to overlook the internal values of the states.
+This can lead to the "noisy TV" problem, where agents fixate on highly novel but meaningless regions, thus trapping them in suboptimal behaviors (Mavor-Parker et al.,, 2022).
+In contrast, the latter leverages high-level heuristics to guide agents in extracting meaningful values from learning experiences, which helps stabilize convergence.
+(Ma et al., 2024b, ; Memarian et al.,, 2021; Trott et al.,, 2019).
+To overcome the limitations of existing RS methods and combine the advantages of exploration-encouraged and inherent value-based rewards, this paper introduces a novel Self-Adaptive Success Rate based reward shaping mechanism (SASR).
 The success rate, defined as the ratio of a state’s presence in successful trajectories to its total occurrences, works as an auxiliary reward distilled from historical experience.
-The success rate assesses a state’s contribution toward successful task completion, which aligns with the agent’s original objectives, offering informative guidance for agent learning.
+This success rate assesses a state’s contribution toward successful task completion, which closely aligns with the agent’s original objectives, offering informative guidance for learning.
 Furthermore, to mitigate overconfidence caused by deterministic success rates, we adopt Beta distributions to model success rates from a probabilistic perspective.
-Beta distributions allow for a self-adaptive evolution of confidence in approximating a state’s success rate.
-This effectively prevents premature certainty and ensures stable convergence to reliable rewards as more data accumulates.
-To derive Beta distributions, we leverage kernel density estimation enhanced by random Fourier features to efficiently estimate success and failure counts, formulating a highly efficient approach.
+Beta distributions enable a self-adaptive evolution of confidence in approximating a state’s success rate, ensuring the system gradually converges to reliable rewards as more data is collected, while preventing premature certainty.
+To derive Beta distributions, we use kernel density estimation (KDE) enhanced by random Fourier features (RFF) to estimate success and failure counts, formulating a highly efficient approach.
 The main contributions of this paper are summarized as follows:
-  1. (i)
-    We propose SASR, an autonomous reward-shaping mechanism without the requirement of expert demonstrations or domain knowledge.
-By extracting a metric from historical experiences that aligns with the agent’s optimization objectives, SASR effectively densifies the environmental rewards.
-  2. (ii)
-    We introduce a novel self-adaptive strategy to achieve a balance between exploration and exploitation.
-Initially, low-confidence Beta distributions provide stochastic rewards that encourage exploration by diversifying the policy optimization directions.
-With the accumulation of experience, high-confidence Beta distributions offer more reliable and precise rewards to enhance exploitation.
-  3. (iii)
-    To derive the Beta distributions in continuous state spaces, we implement kernel density estimation, leveraging random Fourier features.
+  • We propose SASR, an autonomous reward-shaping mechanism without the requirement of domain knowledge.
+By extracting a metric from historical experiences that aligns with the agent’s optimization objectives, SASR effectively augments the environmental rewards.
+  • A novel self-adaptive mechanism is introduced to achieve an efficient exploration-exploitation balance.
+Initially, low-confidence Beta distributions provide stochastic rewards that encourage exploration by diversifying policy optimization.
+As more experience accumulates, high-confidence Beta distributions offer more reliable and precise rewards to enhance exploitation.
+  • To derive the Beta distributions in continuous state spaces, we implement KDE with RFF.
 This formulates a non-parametric and learning-free approach, thereby achieving remarkably low computational complexity.
-  4. (iv)
-    We use sparse-reward and continuous control environments, MuJoCo (Todorov et al., 2012) and simulated robot (de Lazcano et al., 2023), to evaluate the performance of SASR.
-It significantly outperforms several baselines in terms of sample efficiency, learning speed, and convergence stability.
+  • SASR is evaluated on various extremely sparse-reward and continuous control tasks, significantly outperforming several baselines in sample efficiency, learning speed, and convergence stability.
 2 Related Work
-Reward shaping (RS) techniques can be broadly categorized based on the source of learning: either from human knowledge or agent’s own experiences.
-Techniques drawing reward models from human knowledge rely on pre-existing expertise, where methods such as Inverse Reinforcement Learning (IRL) (Arora and Doshi, 2021; Ramachandran and Amir, 2007; Ziebart et al., 2008; Hadfield-Menell et al., 2016) and Inverse Optimal Control (IOC) (Schultheis et al., 2021) aim to derive reward or objective functions from expert demonstrations.
-Another branch within this approach, Imitation Learning (IL) (Ho and Ermon, 2016; Duan et al., 2017; Brown et al., 2020), attempts to clone human behaviors from offline data without explicitly modeling reward structures.
-Following this, transferring the extracted reward models to new tasks has received considerable efforts (Bıyık et al., 2022; Wu et al., 2021; Ellis et al., 2021; Cheng et al., 2021; Adamczyk et al., 2023).
-However, these methods require significant human-driven data and generally show limited adaptability to out-of-distribution (OOD) scenarios.
-Therefore, we focus instead on the autonomous self-learning approaches, which based on the meaning of rewards, can be discussed from two perspectives: intrinsic motivation based and hidden value based rewards.
-Intrinsic motivation based RS explores general heuristics or task-agnostic metrics to encourage exploration in sparse-reward environments.
-Potential-based algorithms define the shaped reward in the form of $\gamma\Phi(s^{\prime})-\Phi(s)$, where $\Phi(\cdot)$ is a potential function, enabling the reward to be nullified in the Bellman equation, thus ensuring the optimal policy remains invariant (Devlin and Kudenko, 2012; Asmuth et al., 2008).
-However, the potential function’s design heavily depends on environmental dynamics, making it more suitable in model-based RL.
-More popular methods incorporate exploration bonuses by rewarding novel states (Devidze et al., 2022; Hong et al., 2018; Eysenbach et al., 2019).
-For instance, count-based strategies maintain visitation counts and assign higher rewards to less frequently visited states.
-To overcome the challenge of state counting in continuous spaces, Tang et al.
-(2017) utilized a hash function to discretize the state space, Bellemare et al.
-(2016) introduced pseudo-counts based on recording probabilities, and Ostrovski et al.
-(2017) used PixelCNN Van den Oord et al.
-(2016).
-Our work adopts a KDE approach to estimate the state counts.
-Additionally, Burda et al.
-(2018) proposed random network distillation to assess state novelty in high-dimensional spaces; Curiosity-driven approaches reward encounters with states that are surprising and out-of-prediction (Sun et al., 2022; Pathak et al., 2017).
-The exploration bonuses branch shows effective exploration enhancement, but only considering novelty while ignoring the underlying values may lead to sub-optimal behaviors.
-A notable example illustrating this limitation is the "noisy TV" problem (Mavor-Parker et al., 2022), where agents may become fixated on irrelevant and dynamic stimuli that yield high novelty but no useful information.
-Hidden value based RS, on the other hand, primarily focuses on task-related signals to reveal the extent to which states contribute to achieving higher environmental rewards and their underlying significance.
-Trott et al.
-(2019) integrated the distance between a state and the target as additional rewards.
-Stadie et al.
-(2020) derived informative reward structures using a Self-Tuning Network (STN) to optimize reward guidance.
-Memarian et al.
-(2021) captured the preferences among different trajectories by ranking them via a trained classifier.
-Zheng et al.
-(2018) minimized the KL-divergence between the learned and the original environmental rewards to synchronize their distributions.
-Mguni et al.
-(2023) used an auxiliary agent that competed against the original agent in a Markov game; Instead, Ma et al.
-(2024) introduced ReLara, a collaborative framework where an assistant reward agent automatically generates rewards to guide the policy agent.
-Moreover, the incorporation of multiple agents or hierarchical structures to share and transfer knowledge through synchronized reward functions is also a popular research direction (Park et al., 2023; Yi et al., 2022; Gupta et al., 2023; Hu et al., 2020).
-While intrinsic motivation rewards directly encourage exploration, hidden value rewards support efficient convergence, Ma et al.
-(2024) combined these two advances by introducing a scheme that evolves the shaped rewards from random perturbations to meaningful metrics.
+Reward shaping (RS) methods can generally be categorized based on the source of learning: either from human knowledge or agent’s own experiences.
+Techniques that derive reward models from human knowledge, such as Inverse Reinforcement Learning (IRL) (Arora and Doshi,, 2021; Ramachandran and Amir,, 2007; Ziebart et al.,, 2008; Hadfield-Menell et al.,, 2016) and Inverse Optimal Control (IOC) (Schultheis et al.,, 2021), aim to extract reward or objective functions from expert demonstrations.
+Following this, transferring the learned reward models to new tasks has received considerable efforts (Bıyık et al.,, 2022; Wu et al.,, 2021; Ellis et al.,, 2021; Cheng et al.,, 2021; Adamczyk et al.,, 2023).
+However, these methods rely heavily on human-generated data and often struggle to adapt to out-of-distribution scenarios.
+Thus, our focus shifts toward autonomous self-learning approaches, which can be further divided into intrinsic motivation-based and inherent value-based rewards based on the nature of the rewards.
+Intrinsic motivation based RS explores general heuristics or task-agnostic metrics to encourage exploration.
+Potential-based algorithms define the shaped reward as $\gamma\Phi(s^{\prime})-\Phi(s)$, where $\Phi(\cdot)$ is a potential function.
+This ensures the reward cancels out in the Bellman equation, preserving the optimal policy (Devlin and Kudenko,, 2012; Asmuth et al.,, 2008; Wiewiora,, 2003).
+However, designing the potential function is highly dependent on the environmental dynamics, making it more applicable to model-based RL.
+More commonly, methods incorporate exploration bonuses to reward novel states (Devidze et al.,, 2022; Badia et al.,, 2020; Hong et al.,, 2018; Eysenbach et al.,, 2019).
+Representatively, count-based strategies track visitation counts and assign higher rewards to less frequently visited states (Lobel et al.,, 2023; Machado et al.,, 2020; Fox et al.,, 2018; Fu et al.,, 2017; Martin et al.,, 2017).
+In continuous spaces, state counting is challenging, so Tang et al., (2017) introduced a hash function to discretize the state space, Bellemare et al., (2016) proposed pseudo-counts based on recording probabilities, and Ostrovski et al., (2017) used PixelCNN Van den Oord et al., (2016).
+Additionally, random network distillation based methods measure state novelty by neural networks (Yang et al.,, 2024; Burda et al.,, 2018), curiosity-driven approaches reward agents for encountering surprising or unpredictable states (Sun et al.,, 2022; Burda et al.,, 2019; Pathak et al.,, 2017).
+Although intrinsic motivation has proven effective in enhancing exploration, only considering novelty while ignoring the inherent values of states may lead to suboptimal policies.
+A notable example illustrating this limitation is the "noisy TV" problem (Mavor-Parker et al.,, 2022), where agents may become attracted to irrelevant, dynamic stimuli that generate high novelty but provide no useful information.
+Inherent value based RS, on the other hand, focuses on task-related signals that highlight how states contribute to achieving higher rewards and their underlying significance.
+For instance, Trott et al., (2019) introduced additional rewards based on the distance between a state and the target; Stadie et al., (2020) derived informative reward structures using a Self-Tuning Network to optimize guidance; Memarian et al., (2021) captured the preferences among different trajectories by ranking them via a trained classifier; Zheng et al., (2018) minimized the KL-divergence between learned and original rewards to align their distributions.
+Mguni et al., (2023) used an auxiliary agent competing against the original agent in a Markov game; Instead, Ma et al., 2024b introduced ReLara, a collaborative framework where an assistant reward agent automatically generates rewards to guide the policy agent.
+Moreover, incorporating multiple agents or hierarchical structures to share and transfer knowledge through synchronized reward functions is another promising research direction (Gupta et al.,, 2023; Ma et al., 2024c, ; Ma et al., 2024a, ; Park et al.,, 2023; Ma et al.,, 2023; Yi et al.,, 2022; Hu et al.,, 2020; Raileanu and Rocktäschel,, 2020).
+While intrinsic motivation rewards guide exploration, inherent value rewards support convergence, Ma et al., 2024b combined these two advances by introducing a scheme that evolves the shaped rewards from random perturbations into meaningful metrics.
 Injecting random rewards during early training relaxes the assumption that only novel states should be rewarded, suggesting instead that any state can be randomly rewarded.
-Building on this concept, our study considers the success rate, which is the probability that a state will contribute to task completion, as the hidden value.
-We sample this metric from evolving Beta distributions that progressively increase their confidence, thus aligning the process to balance the exploration and exploitation adaptively.
-Compared with ReLara, which relies on a black-box agent, our method derives the shaped rewards directly from historical data, effectively reducing the delay of reward convergence.
+Building on this idea, our study considers the success rate as an informative value.
+We sample this metric from evolving Beta distributions that progressively increase in confidence, enabling a self-adaptive exploration-exploitation balance.
+Compared with ReLara, which relies on a black-box agent, our method draws information from historical data, effectively reducing the delay of reward convergence and enhancing explainability.
 3 Preliminaries
@@
 An MDP represented as $\langle S,A,T,R,\gamma\rangle$, involves four main components: $S$ is the state space, $A$ is the action space, $T:S\times A\times S\rightarrow[0,1]$ is the probability of transitioning from one state to another given a specific action, and $R:S\rightarrow\mathbb{R}$ is the reward model.
-The objective in RL is to learn a policy $\pi(a|s)$ that maximizes the expected cumulative rewards $G=\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t})]$, where $\pi(a|s)$ indicates the likelihood of selecting action $a$ in state $s$, and $\gamma$ is the discount factor.
-Model-free RL methods include updating value functions, directly optimizing policies, or employing an actor-critic framework (Sutton and Barto, 2018).
+The objective in RL is to learn a policy $\pi(a|s)$ that maximizes the expected cumulative rewards $G=\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t})]$, where $\pi(a|s)$ indicates the likelihood of selecting action $a$ in state $s$, and $\gamma$ is the discount factor (Sutton and Barto,, 2018).
 Beta Distribution is defined on the interval $[0,1]$, making it ideal for modeling proportions or probabilities.
@@
 4 Methodology
-We propose a Self-Adaptive Success Rate based reward shaping mechanism (SASR) to accelerate RL algorithms in sparse-reward environments.
-Figure 1 illustrates the principles of the SASR mechanism: The diagram consists of two parts representing the early and late stages of learning.
-From the early stage to the late stage, the Beta distributions modeling the estimation of the true success rates evolve from being more stochastic to more deterministic as experiences accumulate.
-This progression closely aligns with the agent’s exploration-exploitation balance.
-The process of Beta distributions’ evolution and generation of shaped rewards from them are detailed in Section 4.1.
+We propose a Self-Adaptive Success Rate based reward shaping mechanism (SASR) to accelerate RL algorithms learning in extremely-sparse-reward environments.
+Figure 1 illustrates the principles of the SASR mechanism: The diagram consists of two parts representing the early and late learning stages.
+As experiences accumulate with learning progresses, the Beta distributions modeling the success rates evolve from being more stochastic to more deterministic.
+This autonomous adaption closely aligns with the agent’s exploration-exploitation balance.
+Section 4.1 introduces how Beta distributions evolve and how shaped rewards are generated from them.
 Additionally, to achieve highly efficient computation, we leverage KDE and RFF to estimate success and failure counts, which are used to derive the corresponding Beta distributions, as detailed in Section 4.2.
@@
 4.1 Self-Adaptive Success Rate Sampling
-We consider the following augmented reward function for SASR:
-$R^{SASR}(s)=R^{E}(s)+\lambda R^{S}(s),$ (3)
-where $R^{E}(s)$ is the environmental reward, $R^{S}(s)$ is the additional shaped reward, and $\lambda$ is a weight factor to control its relative scale.
-When $\lambda\rightarrow 0$, we can recover the original reward model.
+We formulate the augmented reward function in SASR as adding an auxiliary shaped reward $R^{S}(s)$ to the environmental reward $R^{E}(s)$, weighting by a factor $\lambda$:
+$R^{SASR}(s)=R^{E}(s)+\lambda R^{S}(s).$ (3)
 We assign the shaped reward $R^{S}(s_{i})$ of a given state based on its success rate – defined as the ratio of the state’s presence in successful trajectories to its total occurrences.
-This metric provides a meaningful and informative reward from a statistical perspective, as a higher success rate indicates a greater likelihood and contribution for the state to guide the agent towards successful task completion, thus deserving a higher shaped reward.
+This metric provides a meaningful reward from a statistical perspective: a higher success rate, reflected in a higher shaped reward, indicates a greater likelihood that the state will guide the agent toward successful task completion.
 Formally, the success rate based shaped reward $R^{S}(s_{i})$ is given by:
 $R^{S}(s_{i})=f\left(\frac{N_{S}(s_{i})}{N_{S}(s_{i})+N_{F}(s_{i})}\right),$ (4)
-where $N_{S}(s_{i})$ and $N_{F}(s_{i})$ denote the counts of appearances for state $s_{i}$ in historical successful and failed trajectories respectively, and $f(\cdot)$ is a mapping function that adjusts the magnitude of rewards: $f:[0,1]\rightarrow[R_{\min},R_{\max}]$.
-Given the counts $N_{S}(s_{i})$ and $N_{F}(s_{i})$ derived from prior experiences, directly using a deterministic success rate $\frac{N_{S}(s_{i})}{N_{S}(s_{i})+N_{F}(s_{i})}$ may lead to overconfidence in the estimation of the true value.
-To address this, we model the success rate estimation from a probabilistic perspective, inspired by the principles of Thompson sampling, an effective algorithm widely used to balance exploration and exploitation in multi-armed bandit problems (Thompson, 1933; Agrawal and Goyal, 2012).
+where $N_{S}(s_{i})$ and $N_{F}(s_{i})$ denote the counts of state $s_{i}$ appearing in successful and failed historical trajectories, respectively, and $f(\cdot)$ is a mapping function that adjusts the magnitude of rewards: $f:[0,1]\rightarrow[R_{\min},R_{\max}]$.
+Given $N_{S}(s_{i})$ and $N_{F}(s_{i})$, directly using a deterministic success rate $N_{S}(s_{i})/(N_{S}(s_{i})+N_{F}(s_{i}))$ may lead to overconfidence in the estimation of the true value.
+To address this, we model the success rate estimation from a probabilistic perspective, inspired by the principles of Thompson sampling, an effective algorithm widely used to balance exploration and exploitation in multi-armed bandit problems (Thompson,, 1933; Agrawal and Goyal,, 2012).
 We model the success rate of each state as an approximated variable in a Beta distribution, setting the shape parameters $\alpha=N_{S}(s_{i})+1$ and $\beta=N_{F}(s_{i})+1$:
@@
 where the beta function $B(\cdot,\cdot)$ is the normalization factor.
-By sampling from this distribution, we obtain a probabilistic estimate of the state’s success rate.
-This sampled value, $r^{S}_{i}$, is then processed through the mapping function $f(\cdot)$ to produce the shaped reward: $R^{S}(s_{i})=f(r^{S}_{i})$.
+By sampling from this distribution, we obtain a probabilistic estimate of the true success rate.
+This sampled value, $r^{S}_{i}$, is then processed through the scaling function $f(\cdot)$ to produce the shaped reward: $R^{S}(s_{i})=f(r^{S}_{i})$.
 As $N_{S}(s_{i})$ and $N_{F}(s_{i})$ progressively increase throughout the learning process, they influence the shape and sampling variability of the Beta distribution.
@@
   • Encourage Exploration.
-In the early stages of learning, the counts $N_{S}(s_{i})$ and $N_{F}(s_{i})$ are relatively low, resulting in broader Beta distributions and greater variability in sampled values, which injects significant randomness into the shaped rewards.
+In the early learning phases, the counts $N_{S}(s_{i})$ and $N_{F}(s_{i})$ are relatively low, resulting in broader Beta distributions and greater variability in sampled values, which injects significant randomness into the shaped rewards.
 These random rewards optimize the agent in various directions with tiny steps, thus shifting the anchors from which the stochastic policy samples actions.
 The strategy effectively broadens the range of policies and actions, thus encouraging the exploration.
-(Supporting experimental results are provided in Section 5.2.)
+(Supporting experimental results in Section 5.2.)
   • Enhance Exploitation.
-In the late stage of learning, as the counts $N_{S}(s_{i})$ and $N_{F}(s_{i})$ increase, the Beta distribution gradually sharpens, concentrating generated rewards closer to the true success rate.
-The evolution to a certain reward signal with higher confidence highly supports the agent’s exploitation phase, facilitating more efficient convergence towards optimal policies.
+In the late learning phases, as the counts $N_{S}(s_{i})$ and $N_{F}(s_{i})$ increase, the Beta distribution gradually sharpens, concentrating generated rewards closer to the true success rate.
+The evolution to a certain reward signal with higher confidence highly supports the agent’s exploitation, facilitating more efficient convergence towards optimal policies.
   • Consistent Optimization.
-The peak of the Beta distribution, $B(N_{S}(s_{i})+1,N_{F}(s_{i})+1)$, is given by $\frac{N_{S}(s_{i})}{N_{S}(s_{i})+N_{F}(s_{i})}$, directly approximating the success rate.
-Meanwhile, the expected value of the distribution, $\frac{N_{S}(s_{i})+1}{N_{S}(s_{i})+N_{F}(s_{i})+2}$, aligns closely with the calculated success rate.
-This ensures that, despite some randomness, the overall reward remains consistent with policy optimization.
+The peak of the Beta distribution, computed as $N_{S}(s_{i})/(N_{S}(s_{i})+N_{F}(s_{i}))$, directly equals to the success rate.
+Meanwhile, the expectation, $(N_{S}(s_{i})+1)/(N_{S}(s_{i})+N_{F}(s_{i})+2)$, closely approximates the success rate.
+This ensures that, despite stochastic, the overall reward remains consistent with policy optimization.
 4.2 Highly Efficient Beta Distribution Derivation
-In this section, we present the methods to obtain the success and failure counts, $N_{S}(s_{i})$ and $N_{F}(s_{i})$, which are required for deriving the Beta distributions.
+In this section, we present the methods to obtain the success and failure counts, $N_{S}(s_{i})$ and $N_{F}(s_{i})$, for deriving the Beta distributions.
 In tabular environments with low-dimensional, discrete, and finite state spaces, tracking these counts can be straightforwardly accomplished using two tables.
@@
 $\tilde{d}_{X}(s_{i})=\frac{1}{|\mathcal{D}_{X}|}\sum_{j=1}^{|\mathcal{D}_{X}|}{K(s_{i}-s_{j})},\quad X\in\{S,F\},$ (6)
-where $K(\cdot)$ is the kernel function, and $|\mathcal{D}_{X}|$ is the total number of states in the corresponding buffer.
-We select the Gaussian kernel in our implementation.
-The estimated density $\tilde{d}_{X}(s_{i})$ serves as a proxy for the likelihood of encountering state $s_{i}$ under success or failure scenarios, providing a statistically sound basis for estimating $N_{X}(s_{i})$.
-By multiplying $\tilde{d}_{X}(s_{i})$ with the total number of observed states $N$, and representing states as $k$-dimensional vectors, the count $\tilde{N}_{X}(s_{i})$ is estimated as:
-$\tilde{N}_{X}(\bm{s}_{i})=N\times\tilde{d}_{X}(\bm{s}_{i})=\frac{N}{|\mathcal{D}_{X}|}\sum_{j=1}^{|\mathcal{D}_{X}|}{\exp\left(-\frac{||\bm{s}_{i}-\bm{s}_{j}||^{2}}{2\sigma^{2}}\right)},\quad X\in\{S,F\},$ (7)
-where $\sigma$ is a hyperparameter that controls the bandwidth of the Gaussian kernel.
-We further integrate Random Fourier Features (RFF) (Rahimi and Recht, 2007) to reduce computational complexity, as calculating the Gaussian kernel can be expensive, especially in scenarios involving high-dimensional state spaces and large buffers.
+where $K(\cdot)$ is the kernel function and $|\mathcal{D}_{X}|$ is the buffer size.
+We select Gaussian kernel in our implementation.
+The estimated density $\tilde{d}_{X}(s_{i})$ approximates the likelihood of encountering state $s_{i}$ in success or failure scenarios, providing a statistically sound basis for estimating $N_{X}(s_{i})$.
+By multiplying $\tilde{d}_{X}(s_{i})$ by the total number of observed states $N$, the count $\tilde{N}_{X}(s_{i})$ is estimated as:
+$\tilde{N}_{X}(\bm{s}_{i})=N\times\tilde{d}_{X}(\bm{s}_{i})=\frac{N}{|\mathcal{D}_{X}|}\sum_{j=1}^{|\mathcal{D}_{X}|}{\exp\left(-\frac{||\bm{s}_{i}-\bm{s}_{j}||^{2}}{2h^{2}}\right)},\quad X\in\{S,F\},$ (7)
+where $h$ is a hyperparameter, the bandwidth of the Gaussian kernel.
+We further integrate Random Fourier Features (RFF) (Rahimi and Recht,, 2007) to reduce computational complexity, as calculating the Gaussian kernel can be expensive, especially in scenarios involving high-dimensional state spaces and large buffers.
 RFF approximates the kernel function of the original $k$-dimensional states through an inner product of $M$-dimensional randomized features:
@@
 $\bm{w}^{(m)}\sim\mathcal{N}(\bm{0},\sigma^{-2}\bm{I}_{k}),\quad b^{(m)}\sim\text{Uniform}(0,2\pi),\quad m=1,\ldots,M,$ (9)
-where $\bm{I}_{k}$ is the $k\times k$ identity matrix.
-Equation 9 is applied for the Gaussian kernel, while different kernels, such as Laplacian or Cauchy kernels, have different spectral distributions.
-Detailed derivations of the RFF method and the list of spectral distributions are provided in Appendix A.
-Suppose for each iteration, the buffer size of $\mathcal{D}_{X}$ is $D$ and the batch size is $B$, the computational complexity to retrieve the corresponding count $N_{X}$ is $O(MDB)$, representing a linear complexity (detailed derivation in Appendix B).
-RFF transforms the nonlinear kernel computations into vectors’ linear operations, significantly accelerating computation by leveraging the vectorization capabilities of GPUs (Dongarra et al., 2014).
-This demonstrates its higher efficiency compared to other methods that rely on network updates and inferences involving extensive nonlinear computations.
-Moreover, we introduce a hyperparameter, the retention rate $\phi\in(0,1]$, to regulate the volume and diversity of states stored in the buffers.
-Instead of storing all the states in each trajectory, we uniformly retain a specific portion $\phi$.
-This strategy offers two main benefits: (1) as adjacent states in one trajectory tend to be highly similar, especially those closer to the initial state are repetitive and uninformative, retaining a portion of states can skip redundant states and increase sample diversity; (2) maintaining a lower retention rate in the early training stage keeps lower $N_{S}$ and $N_{F}$, resulting in a broader Beta distribution and avoiding premature overconfidence.
-Given the retention rate $\phi$, the space complexity of maintaining two buffers is $O(\phi T|s|)$, where $T$ is the total number of iterations and $|s|$ is the size of a single state.
-To address scalability concerns in more complex or high-dimensional scenarios, we propose leveraging the existing replay buffer of the backbone algorithm, soft actor-critic.
-By augmenting this buffer with a flag array that indicates whether each state is a success or failure, while this introduces additional indexing computations during processing, we can efficiently save space requirements.
+where $\bm{I}_{k}$ is the $k\times k$ identity matrix. Equation 9 is applied for the Gaussian kernel, while different kernels and the detailed derivations of the RFF method are provided in Appendix A.
+4.2.1 Time and Space Complexity of SASR
+Suppose the buffer size of $\mathcal{D}_{X}$ is $D$ and the batch size is $B$ per iteration, the computational complexity to compute the count $N_{X}$ is $O(MDB)$, indicating linear complexity (detailed derivation in Appendix B).
+RFF transforms nonlinear kernel computations into linear vector operations, significantly speeding up computation by leveraging the vectorization capabilities of GPUs (Dongarra et al.,, 2014).
+This demonstrates its higher efficiency compared to methods that rely on network updates and inferences involving extensive nonlinear computations.
+For an empirical comparison of inference time with and without RFF, as well as related discussions, please refer to Appendix C.
+Given the retention rate $\phi$, the space complexity of maintaining two buffers is $O(\phi T|s|)$, where $T$ is the total iterations and $|s|$ is the size of a single state.
+However, storage space can be largely saved by leveraging the existing replay buffer in off-policy backbone RL algorithms, such as SAC (Haarnoja et al., 2018a, ) and TD3 (Fujimoto et al.,, 2018).
+By augmenting the replay buffer with a flag that marks each state as either success or failure, we can efficiently manage space requirements despite some additional indexing computations introduced during processing.
+4.2.2 Implementation Details
+Retention Rate.
+We introduce a hyperparameter, the retention rate $\phi\in(0,1]$, to regulate the volume and diversity of states stored in the buffers.
+Rather than storing all encountered states, we uniformly retain a specific portion of $\phi$.
+The motivation behind this are: (1) adjacent states in one trajectory tend to be highly similar, especially those near the initial state are repetitive and uninformative, retaining a portion of states can skip redundant states and increase sample diversity; (2) using a lower retention rate in the early stage keeps $N_{S}$ and $N_{F}$ lower, resulting in broader Beta distributions and preventing premature overconfidence.
+Defining Success and Failure.
+In tasks where sparse rewards are only given at the end of an episode to indicate task completion, the entire trajectory can be classified as either a success or failure based on the episodic reward.
+For tasks with sparse rewards that do not explicitly indicate task completion, we segment the trajectories by positive reward occurrences.
+Specifically, if a reward is obtained within a pre-defined maximum steps, the corresponding sub-sequence is classified as a success; otherwise, it is considered a failure.
 4.3 The SASR Mechanism for RL agents
+Building upon the SASR reward, we employ the soft actor-critic (SAC) algorithm by Haarnoja et al., 2018a as the foundation for our RL agent.
+Let $Q^{\mathcal{A}_{P}}_{\psi}$ be parameterized Q-network and $\pi^{\mathcal{A}_{P}}_{\theta}$ be parameterized policy network.
+The Q-network can be optimized by the following loss function:
+$\mathcal{L}(\psi)=\Big{(}Q_{\psi}(s_{t},a_{t})-\big{(}r^{E}_{t}+\lambda R^{S}(s_{t})+\gamma Q_{{\psi}^{\prime}}(s_{t+1},a_{t+1})\big{)}\Big{)}^{2},$ (10)
+where $Q_{{\psi}^{\prime}}$ is obtained from a secondary frozen target network to maintain a fixed objective (Mnih et al.,, 2015).
+It is worth noting that the environmental reward $r^{E}_{t}$ is retrieved from the replay buffer.
+Conversely, the shaped reward $R^{S}(s_{t})$ is computed in real-time using the most recently updated $N_{S}(s_{t})$ and $N_{F}(s_{t})$, thereby capturing the latest advancements in learning progress.
+We optimize the policy network by maximizing the expected Q-value and the entropy of the policy $\mathcal{H}\big{(}\pi^{\mathcal{A}_{P}}(\cdot|s_{t})\big{)}$, following Haarnoja et al., 2018b :
+$\mathcal{L}(\theta)=\mathbb{E}_{a_{t}\sim\pi_{\theta}(\cdot|s_{t})}\big{[}-Q_{\psi}(s_{t},a_{t})+\log{\pi_{\theta}(a_{t}|s_{t})}\big{]}.$ (11)
 Algorithm 1 Self-Adaptive Success Rate based Reward Shaping
@@
 17:  end for
-Building upon the SASR reward, we employ the soft actor-critic (SAC) algorithm of Haarnoja et al.
-(2018a) as the foundation for our RL agent.
-Let $\pi^{\mathcal{A}_{P}}_{\theta}$ be parameterized policy network and $Q^{\mathcal{A}_{P}}_{\phi}$ be parameterized Q-network.
-We optimize the Q-function by minimizing the mean squared error (MSE) between the predicted Q-value and the temporal difference (TD) target:
-$\mathcal{L}(\phi)=\Big{(}Q_{\phi}(s_{t},a_{t})-\big{(}r^{E}_{t}+\lambda R^{S}(s_{t})+\gamma Q_{{\phi}^{\prime}}(s_{t+1},a_{t+1})\big{)}\Big{)}^{2},$ (10)
-where $Q_{{\phi}^{\prime}}$ is obtained from a secondary frozen target network to maintain a fixed objective (Mnih et al., 2015).
-It is worth noting that the environmental reward $r^{E}_{t}$ is retrieved from the replay buffer.
-Conversely, the shaped reward $R^{S}(s_{t})$ is computed in real-time using the most recently updated $N_{S}(s_{t})$ and $N_{F}(s_{t})$, thereby capturing the latest advancements in learning progress.
-We optimize the policy network by maximizing the expected Q-value and the entropy of the policy $\mathcal{H}\big{(}\pi^{\mathcal{A}_{P}}(\cdot|s_{t})\big{)}$ (Haarnoja et al., 2018b):
-$\mathcal{L}(\theta)=\mathbb{E}_{a_{t}\sim\pi_{\theta}(\cdot|s_{t})}\big{[}-Q_{\phi}(s_{t},a_{t})+\log{\pi_{\theta}(a_{t}|s_{t})}\big{]}.$ (11)
-The outline of the SASR is summarized in Algorithm 1.
+The flow of the SAC-embedded SASR algorithm is summarized in Algorithm 1.
 We maintain two buffers to store successful and failed states and estimate the success and failure counts from them.
@@
 The shaped rewards, combined with environmental rewards, are used to update the agent’s Q-function and policy.
-In summary, SASR offers two primary advantages: First, it leverages the adaptive properties of Beta distributions to continuously adjust the agent’s confidence levels in understanding the environment, effectively maintaining the balance between exploration and exploitation.
-Second, SASR operates as a non-parametric, learning-free approach, enabling highly efficient reward calculation.
 5 Experiments
-We evaluate SASR on continuous control environments, including seven MuJoCo tasks (Todorov et al., 2012): AntStand, AntSpeed, AntFar, AntVerFar, WalkerKeep, HumanStand, and HumanKeep, two robotic tasks (de Lazcano et al., 2023): RobotReach and RobotPush, as well as one physical simulation task MountainCar (Towers et al., 2023), shown in Figure 2.
-Our environment setup follows the guidelines established in the ReLara paper (Ma et al., 2024), with ReLara also included as a baseline for comparison.
-The tasks provide only sparse rewards, with a reward of $1$ granted upon reaching the final objective within the maximum permitted steps, and 00 for all other states.
-To ensure robust experimental validation, we train multiple instances for each setting using different random seeds and report the average results.
-Additionally, we maintain uniform hyperparameters and network architectures across all experiments, as detailed in Appendix D.
+We evaluate SASR in continuous control environments, following the guidelines established by Ma et al., 2024b , including seven MuJoCo tasks (Todorov et al.,, 2012), two robotic tasks (de Lazcano et al.,, 2023), and one physical simulation task (Towers et al.,, 2023), as shown in Figure 2.
+All tasks provide extremely sparse rewards, with a reward of $1$ granted only upon reaching the final objective within the maximum permitted steps.
+To ensure robust validation, we run $10$ instances for each setting using different random seeds and report the average results.
+We also maintain consistent hyperparameters and network architectures across all experiments, as detailed in Appendix E.
 AntStand
@@
 WalkerKeep
+HumanStand
 HumanKeep
-HumanStand
 RobotReach
@@
 MountainCar
-Figure 2: MuJoCo, robotic and physical simulation tasks with their objectives in our experiments.
+Figure 2: MuJoCo, robotic and physical simulation tasks with their objectives in our experiments. Detailed descriptions of each task are provided in Appendix G.
 5.1 Comparison and Discussion
 Baselines.
-We compare SASR with eight baselines to benchmark its performance: (a) RL with an Assistant Reward Agent (ReLara) (Ma et al., 2024), (b) RL Optimizing Shaping Algorithm (ROSA) (Mguni et al., 2023), (c) Exploration-Guided RS (ExploRS) (Devidze et al., 2022), (d) Count-based static hashing exploration (#Explo) (Tang et al., 2017), (e) Random Network Distillation (RND) (Burda et al., 2018), (f) Soft Actor-Critic (SAC) (Haarnoja et al., 2018a), (g) Twin Delayed Deep Deterministic Policy Gradient (TD3) (Fujimoto et al., 2018), and (h) Proximal Policy Optimization (PPO) (Schulman et al., 2017).
-Algorithms (a) to (e) are all reward shaping methods, involving either exploration bonuses or auxiliary agents to provide additional rewards to tackle the sparse-reward challenge, while algorithms (f) to (h) are advanced RL algorithms.
-For the implementation of these baselines, we employ either the CleanRL library (Huang et al., 2022) or codes provided by the authors.
+We compare SASR with eight baselines to benchmark its performance: (a) RL with an Assistant Reward Agent (ReLara) (Ma et al., 2024b, ), (b) RL Optimizing Shaping Algorithm (ROSA) (Mguni et al.,, 2023), (c) Exploration-Guided RS (ExploRS) (Devidze et al.,, 2022), (d) Count-based static hashing exploration (#Explo) (Tang et al.,, 2017), (e) Random Network Distillation (RND) (Burda et al.,, 2018), (f) Soft Actor-Critic (SAC) (Haarnoja et al., 2018a, ), (g) Twin Delayed DDPG (TD3) (Fujimoto et al.,, 2018), and (h) Proximal Policy Optimization (PPO) (Schulman et al.,, 2017).
+Algorithms (a) to (e) are all reward shaping methods, involving either exploration bonuses or auxiliary agents to shape rewards, while algorithms (f) to (h) are advanced RL algorithms.
 Figure 3 and Table 1 report the average episodic returns and standard errors throughout training.
 Our findings indicate that SASR surpasses the baselines in terms of sample efficiency, learning stability, and convergence speed.
-The primary challenges in these environments are attributed to delayed and sparse rewards alongside long horizons, making exploration essential for timely obtaining successful trajectories.
+The primary challenge in these environments is the extremely sparse reward given after a long horizon, making exploration essential for timely obtaining successful trajectories.
 Although the exploration strategies of algorithms such as ExploRS, #Explo, and RND are designed to reward novel states, effectively expanding the early exploration space with the direct additional target, they continue focusing on discovering novel states, overlooking the implicit values of these states, which makes them fail to return to the final objectives.
@@
 PPO 4.54$\pm$0.04 0.00$\pm$0.00 6.33$\pm$0.24 1.80$\pm$0.11 33.77$\pm$1.11 8.36$\pm$0.03 138.13$\pm$12.64 79.52$\pm$10.80 0.00$\pm$0.00 0.93$\pm$0.00
-SASR, on the other hand, utilizes the success rate as auxiliary information to assess each state’s contribution to task completion.
-This metric is consistent with the agent’s primary objectives, and incorporating this meaningful signal into the temporal difference updating at each step significantly enhances learning efficiency.
-Without this integration, the agent would have to wait until the end of the trajectory to obtain informative signals.
-Furthermore, SASR regulates an efficient exploration strategy by injecting substantial random rewards during the early phases to optimize the agent in multiple directions thereby increasing the likelihood of collecting positive samples.
-This is achieved by sampling from initially wider Beta distributions.
-Importantly, these Beta distributions adaptively yield increasingly confident and accurate success rates as more data is collected, thereby accelerating the agent’s learning.
-These strategies collectively enhance SASR’s efficiency and stability in handling these challenging tasks.
-While ReLara employed a similar exploration mechanism with SASR by rewards perturbations to encourage exploration, it relies on a separate black-box agent, which requires more iterations to converge.
-In contrast, SASR’s success rate sampling is more intuitive and direct, reducing the delay in providing valuable information to the policy agent.
-ReLara’s advantage lies in its integration of the policy agent’s action into the reward construction, as evidenced in the RobotReach task.
-The target point for the robot is randomly selected in each episode, this can be considered as a higher dimension for SASR, because of the changing target.
-In this case, ReLara achieves better results, with its access to action information.
-This can be improved in SASR by incorporating the action as an additional feature in the state vector.
+SASR outperforms the baselines primarily due to its self-adaptive reward evolution mechanism.
+In the early phases, SASR encourages exploration by injecting substantial random rewards to optimize the agent in multiple directions, thereby increasing the likelihood of collecting positive samples.
+Unlike novelty-driven exploration strategies that reward only novel states, SASR relaxes this assumption by rewarding random states, under the belief that all states are worth exploring, and this approach has proven empirically effective.
+Additionally, SASR leverages the success rate as auxiliary information to assess each state’s contribution to task completion, aligning with the agent’s primary objectives.
+As more data is collected, the success rate becomes more accurate, and the shaped reward provides more valuable guidance, significantly enhancing the exploitation and stabilizing the convergence toward optimal policies.
+Cooperatively, these strategies improve SASR’s sample efficiency and convergence stability in challenging tasks.
+While ReLara used a similar exploration mechanism to SASR by perturbing reward functions, it relies on an independent black-box agent, requiring more iterations to converge.
+In contrast, SASR’s success rate sampling is more direct, reducing the delay in converging to valuable information.
+ReLara’s advantage lies in incorporating the policy agent’s actions into reward construction, as seen in the RobotReach task, where the target point is randomly selected in each episode.
+In this case, ReLara outperforms SASR due to access to action information.
+This can also be achieved in SASR by treating actions as additional features in the state vector.
 5.2 Effect of Self-Adaptive Success Rate Sampling
 SASR introduces a novel self-adaptive mechanism that balances exploration and exploitation by maintaining the randomness of the shaped rewards.
-Figure 4 illustrates the generated rewards at five stages of learning in the AntStand task.
-Figure 4 (bottom) represents the learning curve, and Figure 4 (top) illustrates the distribution of the generated rewards over the "height of the ant" feature, one of the dimensions in the state vector.
-Figure 4: Distribution of the shaped rewards over the height of the ant robot in the AntStand task, at different stages of the training process.
-As learning progresses, the evolving shaped rewards exhibit two crucial attributes: the values transition from random to meaningful, and the variance decreases from uncertain to deterministic.
-In the early stages of training, the shaped rewards involve significant randomness due to higher uncertainty.
-Although these random signals contain limited information, they prompt the agent to take tiny optimization steps in diverse directions, effectively shifting the anchors of the policies to various regions.
-This strategy broadens the actions generated by SAC’s stochastic policy, encouraging the agent’s exploration and increasing the diversity of collected samples.
-From another perspective, the perturbations on the rewards can be seen as a relaxation of the assumption that only novel states should be rewarded, instead, randomly reward any state, which also shows its empirical effectiveness.
-In the later stages, the rewards are observed to converge to closely match the height of the robot, a meaningful and intuition-aligned metric, which enhances the agent’s exploitation of its policy.
-This self-adaptive mechanism for balancing exploration and exploitation improves SASR’s ability to gather positive samples earlier, thereby facilitating a faster convergence to the optimal policy.
+To further explore the effect of this mechanism, we use the AntStand task as a case study, analyzing the shaped rewards learned at different stages of training.
+Figure 4 (bottom) shows the learning curve, while Figure 4 (top) illustrates the distributions of generated rewards over the “height of the ant" feature, one dimension in the state vector.
+Figure 4: Distributions of the shaped rewards over the height of the ant robot in the AntStand task at different training stages.
+Red diamonds represent the estimated success rate, while the blue polylines show the actual shaped rewards sampled from the Beta distribution.
+Figure 5: The density of visited states in the MountainCar task for four training periods.
+As learning progresses, the shaped rewards demonstrate two key attributes: the values transition from random to meaningful, and the variance decreases from uncertain to deterministic.
+Besides, although the sampled rewards fluctuate, their center gradually shows a positive linear correlation with the robot’s height.
+In early phases, the shaped rewards contain significant randomness due to higher uncertainty.
+Although these random signals offer limited information, they drive the agent to take tiny optimization steps in diverse directions, effectively shifting the anchors of the policies, thus broadens the actions sampled from SAC’s stochastic policy, encouraging exploration and increasing the diversity of collected samples.
+In later phases, the rewards converge to closely match the height of the robot, a meaningful and intuition-aligned metric, which enhances the agent’s exploitation.
+To further investigate the exploration behavior of SASR, we compare the density of the visited states throughout the training in the MountainCar task with four representative exploration strategies: (1) ReLara, perturbating on both rewards and actions; (2) #Explo, rewarding novel states; (3) SAC, entropy-regularized exploration; and (4) NoisyNet (Fortunato et al.,, 2018), perturbing network weights.
+The density for every 25k steps is shown in Figure 5.
+We observe that SASR progressively covers a wider range of the state space.
+From 50k to 100k steps, SASR reaches positions near the goal, driven by the success rate mechanism.
+In contrast, ReLara covers a similar range to SASR, but is less smooth and takes longer to reach the right side.
+#Explo shows no clear rightward shift, as it focuses on rewarding novelty without considering the inherent value of states.
+SAC’s exploration is relatively narrow, making it easier to trap in local optima.
+NoisyNet shows a narrowing range as network weight perturbations diminish through optimization.
+Overall, SASR demonstrates more effective exploration and collects valuable samples sooner, leading to faster convergence.
 5.3 Ablation Study
-We conduct ablation studies to investigate five key components of SASR: (1) sampling from Beta distributions versus using the deterministic success rates; (2) the retention rate $\phi$, which controls the volume of historical information; (3) the bandwidth $h$ of the kernel density estimation; (4) the number of dimensions $M$ of RFF features; and (5) the weight factor $\lambda$, which adjusts the scales of the shaped rewards.
-We select six representative tasks to report the experimental results, and the quantitative data is provided in Appendix C.
+We conduct ablation studies to investigate key components of SASR. We select six representative tasks to report the experimental results, and the quantitative data is provided in Appendix D.
+(a) The impact of omitting Beta distribution sampling on the performance of SASR.
+(b) Learning performance of SASR with different retention rates $\phi$.
+(c) Learning performance of SASR with different bandwidths $h$ of Gaussian kernel.
+(d) Learning performance of SASR with different RFF feature dimensions $M$.
+(e) Learning performance of SASR with different shaped reward weight factors $\lambda$.
+Figure 6: Ablation study: the impact of key components on the performance of SASR.
 Sampling from Beta distributions.
-We examine a variant of SASR that omits sampling from the Beta distribution and instead directly uses the success rate $N_{S}(s_{i})/(N_{S}(s_{i})+N_{F}(s_{i}))$.
-Comparative results are shown in Figure 5.
-In the early stages of learning, limited experience makes this success rate an inaccurate estimate of the true value.
-Using this overly confident fixed value can mislead the agent.
-Additionally, removing the process of sampling from the Beta distribution disrupts the agent’s exploration encouraged by random rewards, leading to slower convergence.
-The results highlight the importance of sampling from the Beta distribution for effective learning.
-Figure 5: The impact of omitting Beta distribution sampling on the performance of SASR.
-Figure 6: Learning performance of SASR on different retention rates $\phi$.
+(Figure 6(a)) We examine a variant of SASR that omits Beta distribution sampling, instead directly using the success rate $N_{S}(s_{i})/(N_{S}(s_{i})+N_{F}(s_{i}))$.
+In the early stages, limited experience makes this success rate an unreliable estimate, and using this fixed, overly confident value can mislead the agent.
+Furthermore, skipping Beta distribution sampling removes the exploration driven by random rewards, leading to narrower exploration.
+The results highlight the critical role of Beta distribution sampling for effective learning.
 Retention rate $\phi$.
-The retention rate controls the counts’ estimation, subsequently impacting the Beta distribution and sampled rewards.
-We assess various retention rates in Figure 6.
-A high retention rate, such as $\phi=1$, preserving all samples, results in a densely populated and redundant experience pool, causing the Beta distribution to be prematurely overconfident, which can degrade performance.
-Conversely, a lower retention rate $\phi=0.01$ requires more iterations to accumulate sufficient samples, thereby slowing the convergence to meaningful rewards and affecting the agent’s convergence speed.
-The results suggest that carefully balancing the retention rate is crucial, and promising future work could involve adaptive adjustment of the retention rate during the learning process.
+(Figure 6(b)) The retention rate controls the estimation of counts, which affects the confidence of Beta distributions.
+A high retention rate ($\phi=1$) preserves all samples, resulting in a densely populated and redundant state pool, causing the Beta distribution to be prematurely overconfident, which degrades performance.
+Conversely, a low retention rate ($\phi=0.01$) slows convergence by requiring more iterations to gather sufficient samples.
+The results suggest that an appropriate retention rate is important.
 Bandwidth $h$ of Gaussian kernel.
-We evaluate different bandwidths $h$ of the Gaussian kernel for density estimation, as shown in Figure 7.
-Beyond fixed bandwidths, we also test a linearly decreasing bandwidth configuration ($h:0.5\rightarrow 0.1$), which represents progressively increasing confidence in KDE.
-Results indicate that an optimal bandwidth selection is crucial.
-A small bandwidth $h=0.01$ results in farther distances between samples, leading many to have zero estimated density; while a large bandwidth $h=1$ causes the samples indistinguishable due to the overly flat kernel function.
-Both scenarios lead to suboptimal performance.
-The decreasing bandwidth setting does not yield significant improvement and tends to reduce stability due to inconsistent density estimations.
-Figure 7: Learning performance of SASR on different bandwidths $h$ of Gaussian kernel.
-Figure 8: Learning performance of SASR on different RFF feature dimensions $M$.
-Figure 9: Learning performance of SASR on different shaped reward weight factors $\lambda$.
+(Figure 6(c)) The bandwidth $h$ controls the smoothness of the kernel functions.
+Beyond fixed bandwidths, we also test a linearly decreasing configuration ($h:0.5\rightarrow 0.1$), which reflects increasing confidence in KDE.
+Results indicate that a small bandwidth ($h=0.01$) increases the distance between samples, causing many to have zero estimated density, while a large bandwidth ($h=1$) makes samples indistinguishable due to an overly flat kernel function.
+Both cases result in suboptimal performance.
+The decreasing bandwidth setting offers no significant improvement and tends to reduce stability due to inconsistent density estimations.
 RFF feature dimensions $M$.
-Learning performance on different RFF feature dimensions $M$ is shown in Figure 8.
-SASR shows relatively low sensitivity on $M$, provided it is large enough to capture the complexity of the states.
-Results indicate that values like $M=500,1000,2000$ all yield comparable performance levels.
-However, a significantly lower dimension, such as $M=50$, degrades performance, which is also suggested in the original RFF study (Rahimi and Recht, 2007).
+(Figure 6(d)) SASR shows relatively low sensitivity to $M$, provided it is large enough to capture the states’ complexity.
+Results show that values like $M=500,1000,2000$ all yield similar performance levels, while significantly low dimensions, such as $M=50$, degrade performance.
+This is also suggested in the original RFF study (Rahimi and Recht,, 2007).
 Shaped reward weight factor $\lambda$.
-We evaluate the impact of different shaped reward scales on learning performance, as shown in Figure 9.
-We observe that SASR performs better with intermediate values like $\lambda=0.4,0.6,0.8$.
+(Figure 6(e)) Regarding the shaped reward weight factor $\lambda$, we observe that SASR performs better with intermediate values like $\lambda=0.4,0.6,0.8$.
 At $\lambda=0.2$, the minimal shaped reward scale reduces state differentiation, leading to suboptimal performance.
-At $\lambda=1$, aligning the shaped reward scale with the environmental reward introduces excessive uncertainty and randomness, which may overwhelm the feedback disturbances and hinder learning.
-The findings highlight that maintaining a balanced reward scale is important for achieving optimal learning outcomes.
+At $\lambda=1$, aligning the shaped reward scale with the environmental reward introduces excessive uncertainty and randomness, which may overwhelm the feedback and hinder learning.
+The findings highlight that maintaining a balanced reward scale is important for optimal learning outcomes.
 6 Conclusion and Discussion
-We propose SASR in this paper, a self-adaptive reward shaping algorithm based on success rates, to address the sparse-reward challenge.
-SASR achieves an exploration and exploitation trade-off mechanism by generating shaped rewards from evolving Beta distributions.
+In this paper, we propose SASR, a self-adaptive reward shaping algorithm based on success rates, to address the sparse-reward challenge.
+SASR achieves an exploration and exploitation balance mechanism by generating shaped rewards from evolving Beta distributions.
 Experiments demonstrate that this adaptability significantly enhances the agent’s convergence speed.
@@
 This mechanism also provides a sound alternative to traditional count-based RS strategies, adapting effectively to continuous environments.
-Our evaluations confirm the superior performance of SASR in terms of sample efficiency and learning stability, and its robustness across various continuous-control tasks.
-While SASR is designed for environments with sparse rewards, in settings where the environmental rewards are dense, the introduction and maintenance of additional shaped rewards could be unnecessary.
-On the other hand, SASR operates on the premise that sparse rewards indicate task completion to align the success rate notion.
-This correlation is generally straightforward, yet it may not intuitively apply in all contexts.
-Exploring how to extend SASR to more general sparse and even dense reward scenarios presents a promising direction for further research.
-Moreover, the derivation of Beta distributions heavily relies on the samples in the success and failure buffers.
-Currently, our method does not capture the relationships and varying importance of different states within the same trajectory, making the algorithm relatively sensitive to the retention rate, a crucial hyperparameter for achieving promising performance.
-Therefore, developing an adaptive retention rate or better mechanisms for maintaining the two buffers are crucial avenues for future enhancements.
+Our evaluations confirm the superior performance of SASR in terms of sample efficiency and learning stability.
+While SASR is designed for sparse-reward environments, in settings with dense rewards, the introduction and maintenance of additional shaped rewards could be unnecessary.
+Exploring the extension of SASR to dense reward scenarios presents a promising direction for further research.
+Moreover, the derivation of Beta distributions depends heavily on the samples stored in the success and failure buffers.
+Currently, our method does not account for the relationships or varying importance of different states within the same trajectory, making the algorithm sensitive to the retention rate $\phi$.
+Therefore, developing an adaptive retention rate or better mechanisms for managing the buffers are crucial avenues for future improvement.
+7 Acknowledgement
 Appendix A Derivation of Random Fourier Features
-We incorporate Random Fourier Features (RFF) [Rahimi and Recht, 2007] to approximate the kernel functions in the KDE process for the SASR algorithm.
+We incorporate Random Fourier Features (RFF) (Rahimi and Recht,, 2007) to approximate the kernel functions in the KDE process for the SASR algorithm.
 Let the original state be $k$-dimensional, denoted as $\bm{s}\in\mathbb{R}^{k}$, and the kernel function be $k(\bm{s}_{i},\bm{s}_{j})$.
@@
 Kernel Kernel function, $k(\bm{s}_{i}-\bm{s}_{j})$ Spectral density, $p(\bm{w})$
-Gaussian $\exp\left(-\frac{\|\bm{s}_{i}-\bm{s}_{j}\|_{2}^{2}}{2}\right)$ $(2\pi)^{-\frac{M}{2}}\exp\left(-\frac{\|\bm{w}\|_{2}^{2}}{2}\right)$
+Gaussian $\exp\left(-\frac{\|\bm{s}_{i}-\bm{s}_{j}\|_{2}^{2}}{h^{2}}\right)$ $\frac{\sqrt{h}}{2\sqrt{\pi}}\exp\left(-\frac{h\|\bm{w}\|_{2}^{2}}{4}\right)$
 Laplacian $\exp\big{(}-\|\bm{s}_{i}-\bm{s}_{j}\|_{1}\big{)}$ $\prod_{m=1}^{M}\frac{1}{\pi(1+w_{d}^{2})}$
@@
 Since the first two terms $O(DkM)$ and $O(BkM)$ are dominated by the last term $O(MDB)$ when the buffer size and the batch size are large, the overall computational complexity to retrieve the corresponding counts can be approximated as $O(MDB)$.
-Appendix C Supplementary Experimental Results for Ablation Study
+Appendix C Comparison of SASR with and without RFF
+To evaluate the effect of introducing RFF, we compare the training time of SASR with and without RFF, also with the backbone SAC algorithm, the results are shown in Table 3.
+The tests are conducted on the NVIDIA RTX A6000 GPUT.
+The results show that excluding SAC’s inherent optimization time, RFF significantly saves time in the SASR algorithm, while with varying effects across tasks.
+Table 3: Comparison of training time (in hours) for SASR with and without RFF.
+Algorithms AntStand AntFar HumanStand HumanKeep RobotReach RobotPush
+SAC (backbone) 5.87 5.08 4.87 5.67 5.42 6.3
+SASR KDE+RFF 7.15 7.52 6.92 6.20 7.07 8.13
+SASR w/o RFF 8.12 8.72 8.37 6.53 11.12 9.21
+Appendix D Supplementary Experimental Results for Ablation Study
 We provide the detailed quantitative results of the ablation study in this section.
@@
 RobotPush $9.18\pm 1.99$ $118.92\pm 10.79$ $\bm{137.06\pm 12.66}$ $65.51\pm 12.71$ $47.98\pm 9.64$
-Appendix D Network Structures and Hyperparameters
-D.1 Network Structures
+Appendix E Network Structures and Hyperparameters
+E.1 Network Structures
 Figure 7 illustrates the structures of the policy network and Q-network employed in our experiments.
@@
 Figure 7: The structures of policy network and Q-network in our implementation.
-D.2 Hyperparameters
+E.2 Hyperparameters
 We have observed that SASR demonstrated high robustness and was not sensitive to hyperparameter choices. Table 9 shows the set of hyperparameters that we used in all of our experiments.
@@
 burn-in steps $5000$
-Appendix E Compute Resources
+Appendix F Compute Resources
 The experiments in this paper were conducted on a computing cluster, with the detailed hardware configurations listed in Table 10.
@@
 Brand Supermicro 2022
+Appendix G Configurations of Tasks
+We provide the detailed configurations of the tasks in the experiments in this section.
+  • AntStand: The ant robot is trained to stand over a target position. The reward is given if the ant robot reaches the target height. Maximum episode length is 1000 steps.
+  • AntSpeed: The ant robot is trained to move as fast as possible. The reward is given if the ant robot reaches a target speed. Maximum episode length is 1000 steps.
+  • AntFar: The ant robot is trained to reach a target position far from the starting point. The reward is given if the ant robot reaches the target position. Maximum episode length is 1000 steps.
+  • AntVeryFar: The ant robot is trained to reach a target position very far from the starting point.
+The reward is given if the ant robot reaches the target position.
+Maximum episode length is 1000 steps.
+  • WalkerKeep: The walker robot is trained to keep a target height.
+The robot is initialized by standing, and the reward is given if the robot maintains the target height.
+Maximum episode length is 1000 steps.
+  • HumanStand: The human robot is trained to stand over a target position.
+The robot is initialized by lying on the ground, and the reward is given if the robot reaches the target height.
+Maximum episode length is 1000 steps.
+  • HumanKeep: The human robot is trained to keep a target height.
+The robot is initialized by standing, and the reward is given if the robot maintains the target height.
+Maximum episode length is 1000 steps.
+  • RobotReach: The robot arm is trained to reach a target position.
+The target position is randomly generated in the workspace, and the reward is given if the robot reaches the target position.
+Maximum episode length is 500 steps.
+  • RobotPush: The robot arm is trained to push an object to a target position.
+The target position is randomly generated on the table, and the reward is given if the object reaches the target position.
+Maximum episode length is 500 steps.
+  • MountainCar: The car is trained to reach the top of the right hill. The reward is given if the car reaches the top. Maximum episode length is 1000 steps.

### v4

@@
 Abstract
-Reward shaping is a technique in reinforcement learning that addresses the sparse-reward problem by providing more frequent and informative rewards.
-We introduce a self-adaptive and highly efficient reward shaping mechanism that incorporates success rates derived from historical experiences as shaped rewards.
-The success rates are sampled from Beta distributions, which dynamically evolve from uncertain to reliable values as data accumulates.
-Initially, the shaped rewards exhibit more randomness to encourage exploration, while over time, the increasing certainty enhances exploitation, naturally balancing exploration and exploitation.
-Our approach employs Kernel Density Estimation (KDE) combined with Random Fourier Features (RFF) to derive the Beta distributions, providing a computationally efficient, non-parametric, and learning-free solution for high-dimensional continuous state spaces.
-Our method is validated on various tasks with extremely sparse rewards, demonstrating notable improvements in sample efficiency and convergence stability over relevant baselines.
+Reward shaping is a reinforcement learning technique that addresses the sparse-reward problem by providing frequent, informative feedback.
+We propose an efficient self-adaptive reward-shaping mechanism that uses success rates derived from historical experiences as shaped rewards.
+The success rates are sampled from Beta distributions, which evolve from uncertainty to reliability as data accumulates.
+Initially, shaped rewards are stochastic to encourage exploration, gradually becoming more certain to promote exploitation and maintain a natural balance between exploration and exploitation.
+We apply Kernel Density Estimation (KDE) with Random Fourier Features (RFF) to derive Beta distributions, providing a computationally efficient solution for continuous and high-dimensional state spaces.
+Our method, validated on tasks with extremely sparse rewards, improves sample efficiency and convergence stability over relevant baselines.
 1 Introduction
 Environments with extremely sparse rewards present notable challenges for reinforcement learning (RL).
-In such contexts, as the reward model lacks immediate signals, agents receive feedback only after long horizons, making the ability to quickly discover beneficial samples crucial for successful learning (Ladosz et al.,, 2022).
-To address this, a straightforward solution is to reconstruct the reward models by introducing auxiliary signals to assess the agent’s behavior, which has led to the popular technique of Reward Shaping (RS) (Strehl and Littman,, 2008; Gupta et al.,, 2022).
-Inverse reinforcement learning (IRL), which extracts reward functions from human knowledge or expert demonstrations, represents an intuitive approach within this framework Arora and Doshi, (2021).
-However, IRL heavily relies on extensive human input, which can be difficult to obtain especially in complex environments.
+In such contexts, as the reward model lacks immediate signals, agents receive feedback only after long horizons, making the ability to quickly discover beneficial samples crucial for successful learning (Ladosz et al., 2022).
+To address this, a straightforward solution is to reconstruct the reward models by introducing auxiliary signals that assess the agent’s behavior, which has led to the popular technique of Reward Shaping (RS) (Strehl & Littman, 2008; Gupta et al., 2022).
+Inverse reinforcement learning, which extracts reward functions from human knowledge or expert demonstrations, represents an intuitive approach within this framework Arora & Doshi (2021).
+However, it heavily relies on extensive human input, which can be difficult to obtain, especially in complex environments.
 Alternatively, fully autonomous approaches have emerged as an attractive direction.
 Automatically maintained reward shaping can be broadly categorized into two branches: intrinsic motivation-based rewards, which are task-agnostic, and inherent value-based rewards, which are typically task-specific.
-The former mainly introduces exploration bonuses to encourage agents to explore a wider range of states, commonly by rewarding novel or infrequently visited states (Burda et al.,, 2018; Ostrovski et al.,, 2017; Tang et al.,, 2017; Bellemare et al.,, 2016).
+The former mainly introduces exploration bonuses to encourage agents to explore a wider range of states, commonly by rewarding novel or infrequently visited states (Burda et al., 2018; Ostrovski et al., 2017; Tang et al., 2017; Bellemare et al., 2016).
 While these methods effectively enhance exploration, they tend to overlook the internal values of the states.
-This can lead to the "noisy TV" problem, where agents fixate on highly novel but meaningless regions, thus trapping them in suboptimal behaviors (Mavor-Parker et al.,, 2022).
+This can lead to the “noisy TV” problem, where agents fixate on highly novel but meaningless regions, thus trapping them in suboptimal behaviors (Mavor-Parker et al., 2022).
 In contrast, the latter leverages high-level heuristics to guide agents in extracting meaningful values from learning experiences, which helps stabilize convergence.
-(Ma et al., 2024b, ; Memarian et al.,, 2021; Trott et al.,, 2019).
-To overcome the limitations of existing RS methods and combine the advantages of exploration-encouraged and inherent value-based rewards, this paper introduces a novel Self-Adaptive Success Rate based reward shaping mechanism (SASR).
-The success rate, defined as the ratio of a state’s presence in successful trajectories to its total occurrences, works as an auxiliary reward distilled from historical experience.
+However, these methods often struggle in early exploration due to non-directional guidance (Ma et al., 2024a; Memarian et al., 2021; Trott et al., 2019).
+To overcome the limitations of existing RS methods and combine the advantages of exploration-encouraged and inherent value-based rewards, this paper introduces a novel Self-Adaptive Success Rate-based reward shaping mechanism (SASR) [^1] The source code is accessible at: https://github.com/mahaozhe/SASR.
+The success rate, defined as the ratio of a state’s presence in successful trajectories to its total occurrences, serves as an auxiliary reward distilled from historical experience.
 This success rate assesses a state’s contribution toward successful task completion, which closely aligns with the agent’s original objectives, offering informative guidance for learning.
 Furthermore, to mitigate overconfidence caused by deterministic success rates, we adopt Beta distributions to model success rates from a probabilistic perspective.
-Beta distributions enable a self-adaptive evolution of confidence in approximating a state’s success rate, ensuring the system gradually converges to reliable rewards as more data is collected, while preventing premature certainty.
-To derive Beta distributions, we use kernel density estimation (KDE) enhanced by random Fourier features (RFF) to estimate success and failure counts, formulating a highly efficient approach.
+Beta distributions enable a self-adaptive evolution of confidence in approximating a state’s success rate, ensuring the system gradually converges to reliable rewards as more data is collected, while avoiding premature certainty.
+To derive Beta distributions, we use kernel density estimation (KDE) with random Fourier features (RFF) to efficiently estimate success and failure counts.
 The main contributions of this paper are summarized as follows:
-  • We propose SASR, an autonomous reward-shaping mechanism without the requirement of domain knowledge.
-By extracting a metric from historical experiences that aligns with the agent’s optimization objectives, SASR effectively augments the environmental rewards.
-  • A novel self-adaptive mechanism is introduced to achieve an efficient exploration-exploitation balance.
-Initially, low-confidence Beta distributions provide stochastic rewards that encourage exploration by diversifying policy optimization.
-As more experience accumulates, high-confidence Beta distributions offer more reliable and precise rewards to enhance exploitation.
-  • To derive the Beta distributions in continuous state spaces, we implement KDE with RFF.
-This formulates a non-parametric and learning-free approach, thereby achieving remarkably low computational complexity.
-  • SASR is evaluated on various extremely sparse-reward and continuous control tasks, significantly outperforming several baselines in sample efficiency, learning speed, and convergence stability.
+  • We propose SASR, an autonomous reward-shaping mechanism for sparse-reward environments.
+By deriving success rates from historical experiences aligned with the agent’s optimization objectives, SASR effectively augments the environmental rewards.
+  • We introduce a novel self-adaptive mechanism.
+Initially, low-confidence Beta distributions provide uncertain rewards, encouraging exploration by perturbing the reward function and assigning higher rewards to unvisited states.
+As more experience accumulates, high-confidence Beta distributions deliver more reliable and precise rewards to enhance exploitation.
+  • To derive Beta distributions in continuous state spaces, we use KDE with RFF, creating an efficient approach that eliminates the need for additional neural networks or models for learning the auxiliary shaped rewards, thereby achieving remarkably low computational complexity.
+  • SASR is evaluated on various extremely sparse-reward tasks, significantly outperforming several baselines in sample efficiency, learning speed, and convergence stability.
 2 Related Work
-Reward shaping (RS) methods can generally be categorized based on the source of learning: either from human knowledge or agent’s own experiences.
-Techniques that derive reward models from human knowledge, such as Inverse Reinforcement Learning (IRL) (Arora and Doshi,, 2021; Ramachandran and Amir,, 2007; Ziebart et al.,, 2008; Hadfield-Menell et al.,, 2016) and Inverse Optimal Control (IOC) (Schultheis et al.,, 2021), aim to extract reward or objective functions from expert demonstrations.
-Following this, transferring the learned reward models to new tasks has received considerable efforts (Bıyık et al.,, 2022; Wu et al.,, 2021; Ellis et al.,, 2021; Cheng et al.,, 2021; Adamczyk et al.,, 2023).
-However, these methods rely heavily on human-generated data and often struggle to adapt to out-of-distribution scenarios.
-Thus, our focus shifts toward autonomous self-learning approaches, which can be further divided into intrinsic motivation-based and inherent value-based rewards based on the nature of the rewards.
-Intrinsic motivation based RS explores general heuristics or task-agnostic metrics to encourage exploration.
+Reward shaping (RS) methods can generally be categorized based on the source of learning: either from human knowledge or the agent’s own experiences.
+Techniques that derive reward models from human knowledge, such as Inverse Reinforcement Learning (IRL) (Arora & Doshi, 2021; Ramachandran & Amir, 2007; Ziebart et al., 2008; Hadfield-Menell et al., 2016) and Inverse Optimal Control (IOC) (Schultheis et al., 2021; Zhang et al., 2025), aim to extract reward or objective functions from expert demonstrations.
+Subsequently, transferring the learned reward models to new tasks has received considerable efforts (Bıyık et al., 2022; Wu et al., 2021; Ellis et al., 2021; Cheng et al., 2021; Adamczyk et al., 2023; Luo et al., 2024; Lyu et al., 2024).
+However, these methods rely heavily on human-generated data and often struggle adapting out-of-distribution scenarios.
+Thus, our focus shifts toward autonomous self-learning approaches, which can be further divided into intrinsic motivation-based and inherent value-based rewards depending on the nature of the rewards.
+Intrinsic motivation-based RS explores general heuristics or task-agnostic metrics to encourage exploration.
 Potential-based algorithms define the shaped reward as $\gamma\Phi(s^{\prime})-\Phi(s)$, where $\Phi(\cdot)$ is a potential function.
-This ensures the reward cancels out in the Bellman equation, preserving the optimal policy (Devlin and Kudenko,, 2012; Asmuth et al.,, 2008; Wiewiora,, 2003).
-However, designing the potential function is highly dependent on the environmental dynamics, making it more applicable to model-based RL.
-More commonly, methods incorporate exploration bonuses to reward novel states (Devidze et al.,, 2022; Badia et al.,, 2020; Hong et al.,, 2018; Eysenbach et al.,, 2019).
-Representatively, count-based strategies track visitation counts and assign higher rewards to less frequently visited states (Lobel et al.,, 2023; Machado et al.,, 2020; Fox et al.,, 2018; Fu et al.,, 2017; Martin et al.,, 2017).
-In continuous spaces, state counting is challenging, so Tang et al., (2017) introduced a hash function to discretize the state space, Bellemare et al., (2016) proposed pseudo-counts based on recording probabilities, and Ostrovski et al., (2017) used PixelCNN Van den Oord et al., (2016).
-Additionally, random network distillation based methods measure state novelty by neural networks (Yang et al.,, 2024; Burda et al.,, 2018), curiosity-driven approaches reward agents for encountering surprising or unpredictable states (Sun et al.,, 2022; Burda et al.,, 2019; Pathak et al.,, 2017).
-Although intrinsic motivation has proven effective in enhancing exploration, only considering novelty while ignoring the inherent values of states may lead to suboptimal policies.
-A notable example illustrating this limitation is the "noisy TV" problem (Mavor-Parker et al.,, 2022), where agents may become attracted to irrelevant, dynamic stimuli that generate high novelty but provide no useful information.
+This ensures the reward cancels out in the Bellman equation, preserving the optimal policy (Devlin & Kudenko, 2012; Asmuth et al., 2008; Wiewiora, 2003).
+However, designing the potential function highly depends on the environmental dynamics, making it more applicable to model-based RL.
+More commonly, methods incorporate exploration bonuses to reward novel states (Mahankali et al., 2024; Liu et al., 2025; Devidze et al., 2022; Badia et al., 2020; Hong et al., 2018; Eysenbach et al., 2019).
+Count-based strategies, for instance, track visitation counts and assign higher rewards to less frequently visited states (Lobel et al., 2023; Machado et al., 2020; Fox et al., 2018; Fu et al., 2017; Martin et al., 2017).
+In continuous spaces, state counting is challenging, so Tang et al.
+(2017) introduced a hash function to discretize the state space, Bellemare et al.
+(2016) proposed pseudo-counts based on recording probabilities, and Ostrovski et al.
+(2017) used PixelCNN Van den Oord et al.
+(2016) to simulate density.
+Additionally, random network distillation-based methods measure state novelty by neural networks (Yang et al., 2024b; Liu et al., 2024; Burda et al., 2018), while curiosity-driven approaches reward agents for encountering surprising or unpredictable states (Yang et al., 2024a; Sun et al., 2022; Burda et al., 2019; Pathak et al., 2017; Zhang et al., 2023).
+Although intrinsic motivation has proven effective in enhancing exploration, only considering novelty while ignoring the inherent values of states can lead to suboptimal policies.
 Inherent value based RS, on the other hand, focuses on task-related signals that highlight how states contribute to achieving higher rewards and their underlying significance.
-For instance, Trott et al., (2019) introduced additional rewards based on the distance between a state and the target; Stadie et al., (2020) derived informative reward structures using a Self-Tuning Network to optimize guidance; Memarian et al., (2021) captured the preferences among different trajectories by ranking them via a trained classifier; Zheng et al., (2018) minimized the KL-divergence between learned and original rewards to align their distributions.
-Mguni et al., (2023) used an auxiliary agent competing against the original agent in a Markov game; Instead, Ma et al., 2024b introduced ReLara, a collaborative framework where an assistant reward agent automatically generates rewards to guide the policy agent.
-Moreover, incorporating multiple agents or hierarchical structures to share and transfer knowledge through synchronized reward functions is another promising research direction (Gupta et al.,, 2023; Ma et al., 2024c, ; Ma et al., 2024a, ; Park et al.,, 2023; Ma et al.,, 2023; Yi et al.,, 2022; Hu et al.,, 2020; Raileanu and Rocktäschel,, 2020).
-While intrinsic motivation rewards guide exploration, inherent value rewards support convergence, Ma et al., 2024b combined these two advances by introducing a scheme that evolves the shaped rewards from random perturbations into meaningful metrics.
-Injecting random rewards during early training relaxes the assumption that only novel states should be rewarded, suggesting instead that any state can be randomly rewarded.
-Building on this idea, our study considers the success rate as an informative value.
-We sample this metric from evolving Beta distributions that progressively increase in confidence, enabling a self-adaptive exploration-exploitation balance.
-Compared with ReLara, which relies on a black-box agent, our method draws information from historical data, effectively reducing the delay of reward convergence and enhancing explainability.
+For instance, Trott et al.
+(2019) introduced additional rewards based on the distance between a state and the target; Stadie et al.
+(2020) derived informative reward structures using a Self-Tuning Network to optimize guidance; Memarian et al.
+(2021) captured the preferences among different trajectories by ranking them via a trained classifier; Zheng et al.
+(2018) minimized the KL-divergence between learned and original rewards to align their distributions.
+Mguni et al.
+(2023) used an auxiliary agent competing against the original agent in a Markov game; Ma et al.
+(2024a) introduced ReLara, a collaborative framework where a reward agent automatically generates rewards to guide the policy agent.
+Moreover, incorporating multiple agents or hierarchical structures to share and transfer knowledge through synchronized reward functions is another promising research direction (Park et al., 2023; Ma et al., 2024b; Gupta et al., 2023; Hu et al., 2020; Raileanu & Rocktäschel, 2020; Yi et al., 2022).
 3 Preliminaries
@@
 An MDP represented as $\langle S,A,T,R,\gamma\rangle$, involves four main components: $S$ is the state space, $A$ is the action space, $T:S\times A\times S\rightarrow[0,1]$ is the probability of transitioning from one state to another given a specific action, and $R:S\rightarrow\mathbb{R}$ is the reward model.
-The objective in RL is to learn a policy $\pi(a|s)$ that maximizes the expected cumulative rewards $G=\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t})]$, where $\pi(a|s)$ indicates the likelihood of selecting action $a$ in state $s$, and $\gamma$ is the discount factor (Sutton and Barto,, 2018).
+The objective in RL is to learn a policy $\pi(a|s)$ that maximizes the expected cumulative rewards $G=\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t})]$, where $\pi(a|s)$ indicates the likelihood of selecting action $a$ in state $s$, and $\gamma$ is the discount factor (Sutton & Barto, 2018).
 Beta Distribution is defined on the interval $[0,1]$, making it ideal for modeling proportions or probabilities.
-It is parameterized by $\alpha$ and $\beta$, which are interpreted as prior counts of successes and failures of a binary outcome.
+It is parameterized by $\alpha$ and $\beta$, which represent prior counts of successes and failures of a binary outcome.
 The probability density function of a Beta-distributed variable $X$ is:
 $f(x;\alpha,\beta)=\frac{1}{B(\alpha,\beta)}x^{\alpha-1}(1-x)^{\beta-1},$ (1)
-where $B(\alpha,\beta)$ is a normalization factor.
-The key characteristic of Beta distribution is its adaptability: as more data accumulates, the values of $\alpha$ and $\beta$ increase, the distribution’s shape becomes narrower, thereby increasing confidence in the estimated probabilities.
-This feature is particularly beneficial in adaptive online learning, aligning closely with our objective of balancing exploration and exploitation.
-Kernel Density Estimation (KDE) is a non-parametric approach to approximate the probability density function of a random variable from data samples.
+where $B(\alpha,\beta)$ is the beta function.
+The key attribute of Beta distribution is its adaptability: as more data accumulates, the values of $\alpha$ and $\beta$ increase, narrowing the distribution’s shape and increasing confidence in the estimated probabilities.
+This feature is particularly useful in adaptive online learning, aligning with our objective of balancing exploration and exploitation.
+Kernel Density Estimation (KDE) is a non-parametric method for approximating the probability density function of a random variable from data samples.
 Given $n$ data points $\{x_{i}\}_{i=1}^{n}$, KDE smooths these points to approximate the density function as follows:
 $\hat{d}(x)=\frac{1}{nh}\sum_{i=1}^{n}K\left(\frac{x-x_{i}}{h}\right),$ (2)
-where $h$ is the bandwidth, $K(\cdot)$ is a kernel function such as Gaussian kernel, Laplacian kernel, or Cauchy kernel.
-KDE is particularly useful in scenarios where the actual distribution is not well-defined or complex, such as continuous state spaces in our RL environments.
+where $h$ is the bandwidth, and $K(\cdot)$ is a kernel function such as Gaussian kernel, Laplacian kernel, or Cauchy kernel.
+KDE is particularly useful in scenarios where the actual distribution is complex or poorly defined, such as in continuous state spaces in RL environments.
 4 Methodology
-We propose a Self-Adaptive Success Rate based reward shaping mechanism (SASR) to accelerate RL algorithms learning in extremely-sparse-reward environments.
+We propose a Self-Adaptive Success Rate based reward shaping mechanism (SASR) to accelerate RL algorithms in extremely-sparse-reward environments.
 Figure 1 illustrates the principles of the SASR mechanism: The diagram consists of two parts representing the early and late learning stages.
-As experiences accumulate with learning progresses, the Beta distributions modeling the success rates evolve from being more stochastic to more deterministic.
+As experiences accumulate with learning progresses, the Beta distributions modeling the success rates evolve from being stochastic to deterministic.
 This autonomous adaption closely aligns with the agent’s exploration-exploitation balance.
 Section 4.1 introduces how Beta distributions evolve and how shaped rewards are generated from them.
-Additionally, to achieve highly efficient computation, we leverage KDE and RFF to estimate success and failure counts, which are used to derive the corresponding Beta distributions, as detailed in Section 4.2.
+Additionally, to achieve highly efficient computation, we use KDE and RFF to estimate success and failure counts, which are used to derive the corresponding Beta distributions, as detailed in Section 4.2.
 Lastly, Section 4.3 presents the integration of SASR into the RL agent and the overall algorithmic flow.
@@
 $R^{S}(s_{i})=f\left(\frac{N_{S}(s_{i})}{N_{S}(s_{i})+N_{F}(s_{i})}\right),$ (4)
-where $N_{S}(s_{i})$ and $N_{F}(s_{i})$ denote the counts of state $s_{i}$ appearing in successful and failed historical trajectories, respectively, and $f(\cdot)$ is a mapping function that adjusts the magnitude of rewards: $f:[0,1]\rightarrow[R_{\min},R_{\max}]$.
-Given $N_{S}(s_{i})$ and $N_{F}(s_{i})$, directly using a deterministic success rate $N_{S}(s_{i})/(N_{S}(s_{i})+N_{F}(s_{i}))$ may lead to overconfidence in the estimation of the true value.
-To address this, we model the success rate estimation from a probabilistic perspective, inspired by the principles of Thompson sampling, an effective algorithm widely used to balance exploration and exploitation in multi-armed bandit problems (Thompson,, 1933; Agrawal and Goyal,, 2012).
-We model the success rate of each state as an approximated variable in a Beta distribution, setting the shape parameters $\alpha=N_{S}(s_{i})+1$ and $\beta=N_{F}(s_{i})+1$:
+where $N_{S}(s_{i})$ and $N_{F}(s_{i})$ denote the counts of state $s_{i}$ appearing in successful and failed historical trajectories, respectively.
+To enhance scalability and adaptability, $f(\cdot)$ is a linear scaling function that maps the original success rate from $[0,1]$ to a desired scale $[R_{min},R_{max}]$, making the magnitude of the shaped rewards more flexible, i.e., $f(x)=R_{min}+x\cdot(R_{max}-R_{min})$.
+Given $N_{S}(s_{i})$ and $N_{F}(s_{i})$, directly using a deterministic success rate may lead to overconfidence in the estimation of the true value.
+To address this, inspired by the principles of Thompson sampling (Thompson, 1933; Agrawal & Goyal, 2012), we adopt a probabilistic perspective for success rate estimation.
+Specifically, the success rate of each state is approximated as a variable in a Beta distribution, with shape parameters set as $\alpha=N_{S}(s_{i})+1$ and $\beta=N_{F}(s_{i})+1$:
 $r^{S}_{i}\sim\text{Beta}(r;\alpha,\beta)=\frac{1}{B(N_{S}(s_{i})+1,N_{F}(s_{i})+1)}r^{N_{S}(s_{i})}(1-r)^{N_{F}(s_{i})},$ (5)
@@
   • Encourage Exploration.
-In the early learning phases, the counts $N_{S}(s_{i})$ and $N_{F}(s_{i})$ are relatively low, resulting in broader Beta distributions and greater variability in sampled values, which injects significant randomness into the shaped rewards.
-These random rewards optimize the agent in various directions with tiny steps, thus shifting the anchors from which the stochastic policy samples actions.
-The strategy effectively broadens the range of policies and actions, thus encouraging the exploration.
-(Supporting experimental results in Section 5.2.)
+In the early phases, lower counts of $N_{S}(s_{i})$ and $N_{F}(s_{i})$ result in higher-variance Beta distributions, making the sampled rewards more stochastic.
+This acts as a noisy perturbation of the reward function.
+Given that the environmental rewards are mostly zero, this perturbation optimizes the agent in diverse directions through small adjustments, shifting the anchors from which the stochastic policy samples actions.
+Meanwhile, early-visited states are likely to fail, leading to a decrease in their success rates, while unvisited states retain the initial Beta distribution $Beta(1,1)$, receiving relatively higher rewards.
+This mechanism drives the agent to explore novel regions, aligning with the principles of intrinsic motivation.
   • Enhance Exploitation.
-In the late learning phases, as the counts $N_{S}(s_{i})$ and $N_{F}(s_{i})$ increase, the Beta distribution gradually sharpens, concentrating generated rewards closer to the true success rate.
-The evolution to a certain reward signal with higher confidence highly supports the agent’s exploitation, facilitating more efficient convergence towards optimal policies.
+In the later phases, as the counts $N_{S}(s_{i})$ and $N_{F}(s_{i})$ increase, the Beta distribution gradually sharpens, concentrating generated rewards around the true success rate.
+The more certain reward signals with higher confidence highly support the agent’s exploitation, facilitating faster convergence towards optimal policies.
   • Consistent Optimization.
-The peak of the Beta distribution, computed as $N_{S}(s_{i})/(N_{S}(s_{i})+N_{F}(s_{i}))$, directly equals to the success rate.
+The peak of the Beta distribution, given by $N_{S}(s_{i})/(N_{S}(s_{i})+N_{F}(s_{i}))$, equals the success rate.
 Meanwhile, the expectation, $(N_{S}(s_{i})+1)/(N_{S}(s_{i})+N_{F}(s_{i})+2)$, closely approximates the success rate.
-This ensures that, despite stochastic, the overall reward remains consistent with policy optimization.
+This ensures that, despite stochasticity, the overall reward remains consistent with policy optimization.
 4.2 Highly Efficient Beta Distribution Derivation
-In this section, we present the methods to obtain the success and failure counts, $N_{S}(s_{i})$ and $N_{F}(s_{i})$, for deriving the Beta distributions.
-In tabular environments with low-dimensional, discrete, and finite state spaces, tracking these counts can be straightforwardly accomplished using two tables.
-However, most real-world environments involve high-dimensional, continuous, and infinite state spaces, making it impractical to maintain exact counts for every state.
-To address this challenge, we use Kernel Density Estimation (KDE) to approximate the densities of successes and failures from accumulated experience.
+In this section, we present how the success and failure counts, $N_{S}(s_{i})$ and $N_{F}(s_{i})$, are derived for the Beta distributions.
+To efficiently estimate these counts in high-dimensional, continuous, or infinite state spaces, we use Kernel Density Estimation (KDE) to approximate the densities of successes and failures from accumulated experience.
 Specifically, we maintain two buffers, $\mathcal{D}_{S}$ and $\mathcal{D}_{F}$, to store the states in successful and failed trajectories, respectively.
@@
 By multiplying $\tilde{d}_{X}(s_{i})$ by the total number of observed states $N$, the count $\tilde{N}_{X}(s_{i})$ is estimated as:
-$\tilde{N}_{X}(\bm{s}_{i})=N\times\tilde{d}_{X}(\bm{s}_{i})=\frac{N}{|\mathcal{D}_{X}|}\sum_{j=1}^{|\mathcal{D}_{X}|}{\exp\left(-\frac{||\bm{s}_{i}-\bm{s}_{j}||^{2}}{2h^{2}}\right)},\quad X\in\{S,F\},$ (7)
-where $h$ is a hyperparameter, the bandwidth of the Gaussian kernel.
-We further integrate Random Fourier Features (RFF) (Rahimi and Recht,, 2007) to reduce computational complexity, as calculating the Gaussian kernel can be expensive, especially in scenarios involving high-dimensional state spaces and large buffers.
+$\tilde{N}_{X}(s_{i})=N\times\tilde{d}_{X}(s_{i})=\frac{N}{|\mathcal{D}_{X}|}\sum_{j=1}^{|\mathcal{D}_{X}|}{\exp\left(-\frac{||s_{i}-s_{j}||^{2}}{2h^{2}}\right)},\quad X\in\{S,F\},$ (7)
+where hyperparameter $h$ is the bandwidth of the Gaussian kernel.
+We further integrate Random Fourier Features (RFF) (Rahimi & Recht, 2007) to reduce computational complexity, as calculating the Gaussian kernel can be expensive, especially in scenarios involving high-dimensional state spaces and large buffers.
 RFF approximates the kernel function of the original $k$-dimensional states through an inner product of $M$-dimensional randomized features:
-$K(\bm{s}_{i},\bm{s}_{j})\approx z(\bm{s}_{i})^{T}z(\bm{s}_{j}),\quad z(\bm{s})=\sqrt{\frac{2}{M}}\cos(\bm{W}^{T}\bm{s}+\bm{b}),$ (8)
-where $z(\cdot)$ is the RFF mapping function with $\bm{W}=\big{[}\bm{w}^{(1)},\ldots,\bm{w}^{(M)}\big{]}\in\mathbb{R}^{k\times M}$ and $\bm{b}=\big{[}b^{(1)},\ldots,b^{(M)}\big{]}^{T}\in\mathbb{R}^{M}$ randomly sampled from the following spectral distributions:
+$K(s_{i},s_{j})\approx z(s_{i})^{T}z(s_{j}),\quad z(s)=\sqrt{\frac{2}{M}}\cos(\bm{W}^{T}s+\bm{b}),$ (8)
+where $z(\cdot)$ is the RFF mapping function with $\bm{W}=\big{[}\bm{w}^{(1)},\ldots,\bm{w}^{(M)}\big{]}\in\mathbb{R}^{k\times M}$, and $\bm{b}=\big{[}b^{(1)},\ldots,b^{(M)}\big{]}^{T}\in\mathbb{R}^{M}$ is randomly sampled from the following distributions:
 $\bm{w}^{(m)}\sim\mathcal{N}(\bm{0},\sigma^{-2}\bm{I}_{k}),\quad b^{(m)}\sim\text{Uniform}(0,2\pi),\quad m=1,\ldots,M,$ (9)
-where $\bm{I}_{k}$ is the $k\times k$ identity matrix. Equation 9 is applied for the Gaussian kernel, while different kernels and the detailed derivations of the RFF method are provided in Appendix A.
-4.2.1 Time and Space Complexity of SASR
-Suppose the buffer size of $\mathcal{D}_{X}$ is $D$ and the batch size is $B$ per iteration, the computational complexity to compute the count $N_{X}$ is $O(MDB)$, indicating linear complexity (detailed derivation in Appendix B).
-RFF transforms nonlinear kernel computations into linear vector operations, significantly speeding up computation by leveraging the vectorization capabilities of GPUs (Dongarra et al.,, 2014).
-This demonstrates its higher efficiency compared to methods that rely on network updates and inferences involving extensive nonlinear computations.
-For an empirical comparison of inference time with and without RFF, as well as related discussions, please refer to Appendix C.
-Given the retention rate $\phi$, the space complexity of maintaining two buffers is $O(\phi T|s|)$, where $T$ is the total iterations and $|s|$ is the size of a single state.
-However, storage space can be largely saved by leveraging the existing replay buffer in off-policy backbone RL algorithms, such as SAC (Haarnoja et al., 2018a, ) and TD3 (Fujimoto et al.,, 2018).
-By augmenting the replay buffer with a flag that marks each state as either success or failure, we can efficiently manage space requirements despite some additional indexing computations introduced during processing.
+where $\bm{I}_{k}$ is the $k\times k$ identity matrix.
+Equation 9 is applied for the Gaussian kernel, while different kernels and the detailed derivations of the RFF method are provided in Appendix A.1.
 4.2.1 Implementation Details
@@
 Rather than storing all encountered states, we uniformly retain a specific portion of $\phi$.
-The motivation behind this are: (1) adjacent states in one trajectory tend to be highly similar, especially those near the initial state are repetitive and uninformative, retaining a portion of states can skip redundant states and increase sample diversity; (2) using a lower retention rate in the early stage keeps $N_{S}$ and $N_{F}$ lower, resulting in broader Beta distributions and preventing premature overconfidence.
+The motivations behind this are: (1) adjacent states in one trajectory tend to be highly similar, especially those near the initial state are repetitive and uninformative, retaining a portion of states can skip redundant states and increase sample diversity; (2) using a lower retention rate in the early stage keeps $N_{S}$ and $N_{F}$ lower, resulting in broader Beta distributions and preventing premature overconfidence.
 Defining Success and Failure.
@@
 Specifically, if a reward is obtained within a pre-defined maximum steps, the corresponding sub-sequence is classified as a success; otherwise, it is considered a failure.
+4.2.2 Time and Space Complexity of SASR
+Suppose the buffer size of $\mathcal{D}_{X}$ is $D$ and the batch size is $B$ per iteration, the computational complexity to compute the count $N_{X}$ is $O(MDB)$, indicating linear complexity (detailed in Appendix A.2).
+RFF converts nonlinear kernel computations into linear vector operations, significantly speeding up computation by leveraging the vectorization capabilities of GPUs (Dongarra et al., 2014).
+This highlights its superior efficiency compared to methods that rely on network updates and inferences involving extensive nonlinear computations.
+Given the retention rate $\phi$, the space complexity of maintaining two buffers is $O(\phi T|s|)$, where $T$ is the total iterations and $|s|$ is the size of a single state.
+Moreover, storage space is significantly reduced by leveraging the existing replay buffer in off-policy RL algorithms, like SAC (Haarnoja et al., 2018a) and TD3 (Fujimoto et al., 2018).
+Specifically, we augment the replay buffer with a flag that marks each state as either a success ($\text{flag}=1$) or failure ($\text{flag}=0$).
+This allows for the efficient management of space requirements with minimal overhead from indexing.
+For supporting experimental results in time and space complexity, please refer to Appendix A.3.
 4.3 The SASR Mechanism for RL agents
-Building upon the SASR reward, we employ the soft actor-critic (SAC) algorithm by Haarnoja et al., 2018a as the foundation for our RL agent.
-Let $Q^{\mathcal{A}_{P}}_{\psi}$ be parameterized Q-network and $\pi^{\mathcal{A}_{P}}_{\theta}$ be parameterized policy network.
-The Q-network can be optimized by the following loss function:
-$\mathcal{L}(\psi)=\Big{(}Q_{\psi}(s_{t},a_{t})-\big{(}r^{E}_{t}+\lambda R^{S}(s_{t})+\gamma Q_{{\psi}^{\prime}}(s_{t+1},a_{t+1})\big{)}\Big{)}^{2},$ (10)
-where $Q_{{\psi}^{\prime}}$ is obtained from a secondary frozen target network to maintain a fixed objective (Mnih et al.,, 2015).
-It is worth noting that the environmental reward $r^{E}_{t}$ is retrieved from the replay buffer.
-Conversely, the shaped reward $R^{S}(s_{t})$ is computed in real-time using the most recently updated $N_{S}(s_{t})$ and $N_{F}(s_{t})$, thereby capturing the latest advancements in learning progress.
-We optimize the policy network by maximizing the expected Q-value and the entropy of the policy $\mathcal{H}\big{(}\pi^{\mathcal{A}_{P}}(\cdot|s_{t})\big{)}$, following Haarnoja et al., 2018b :
-$\mathcal{L}(\theta)=\mathbb{E}_{a_{t}\sim\pi_{\theta}(\cdot|s_{t})}\big{[}-Q_{\psi}(s_{t},a_{t})+\log{\pi_{\theta}(a_{t}|s_{t})}\big{]}.$ (11)
 Algorithm 1 Self-Adaptive Success Rate based Reward Shaping
@@
 11:     $\{(s_{t},a_{t},r^{E}_{t},s_{t+1})_{i}\}\sim\mathcal{D}$triangleright sample a batch of transitions from the replay buffer
-12:     $\tilde{N}_{S}=z(\bm{s}_{t})^{T}z(\bm{\mathcal{D}_{S}})$triangleright estimate the success counts
-13:     $\tilde{N}_{F}=z(\bm{s}_{t})^{T}z(\bm{\mathcal{D}_{F}})$triangleright estimate the failure counts
+12:     $\tilde{N}_{S}=z(s_{t})^{T}z(\mathcal{D}_{S})$triangleright estimate the success counts
+13:     $\tilde{N}_{F}=z(s_{t})^{T}z(\mathcal{D}_{F})$triangleright estimate the failure counts
 14:     $r^{S}_{t}\sim\text{Beta}(r;\tilde{N}_{S}+1,\tilde{N}_{F}+1)$triangleright sample the success rate from the Beta distribution
@@
 17:  end for
+Building upon the SASR reward, for demonstration, we employ the soft actor-critic (SAC) algorithm by Haarnoja et al.
+(2018a) as the foundational agent.
+Let $Q_{\psi}$ be the parameterized Q-network and $\pi_{\theta}$ be the parameterized policy network.
+The Q-network is optimized by the following loss function:
+$\mathcal{L}(\psi)=\Big{(}Q_{\psi}(s_{t},a_{t})-\big{(}r^{E}_{t}+\lambda R^{S}(s_{t})+\gamma Q_{{\psi}^{\prime}}(s_{t+1},a_{t+1})\big{)}\Big{)}^{2},$ (10)
+where $Q_{{\psi}^{\prime}}$ is obtained from a secondary frozen target network to maintain a fixed objective (Mnih et al., 2015).
+Notably, the environmental reward $r^{E}_{t}$ is retrieved from the replay buffer, conversely, the shaped reward $R^{S}(s_{t})$ is computed in real-time using the most recently updated $N_{S}(s_{t})$ and $N_{F}(s_{t})$, ensuring it reflects the latest learning progress.
+We optimize the policy network by maximizing the expected Q-value and the policy entropy $\mathcal{H}\big{(}\pi_{\theta}(\cdot|s_{t})\big{)}$, following Haarnoja et al. (2018b):
+$\mathcal{L}(\theta)=\mathbb{E}_{a_{t}\sim\pi_{\theta}(\cdot|s_{t})}\big{[}-Q_{\psi}(s_{t},a_{t})+\log{\pi_{\theta}(a_{t}|s_{t})}\big{]}.$ (11)
 The flow of the SAC-embedded SASR algorithm is summarized in Algorithm 1.
-We maintain two buffers to store successful and failed states and estimate the success and failure counts from them.
-Subsequently, shaped rewards are computed by sampling from Beta distributions followed by a scale transformation.
-The shaped rewards, combined with environmental rewards, are used to update the agent’s Q-function and policy.
 5 Experiments
-We evaluate SASR in continuous control environments, following the guidelines established by Ma et al., 2024b , including seven MuJoCo tasks (Todorov et al.,, 2012), two robotic tasks (de Lazcano et al.,, 2023), and one physical simulation task (Towers et al.,, 2023), as shown in Figure 2.
+We evaluate SASR in high-dimensional environments, including four MuJoCo tasks (Todorov et al., 2012), four robotic tasks (de Lazcano et al., 2023), five Atari games, including the well-known Montezuma’s Revenge Bellemare et al.
+(2013), and a physical simulation task (Towers et al., 2023), as shown in Figure 2.
 All tasks provide extremely sparse rewards, with a reward of $1$ granted only upon reaching the final objective within the maximum permitted steps.
-To ensure robust validation, we run $10$ instances for each setting using different random seeds and report the average results.
-We also maintain consistent hyperparameters and network architectures across all experiments, as detailed in Appendix E.
-AntStand
-AntSpeed
-AntFar
-AntVeryFar
-WalkerKeep
-HumanStand
-HumanKeep
-RobotReach
-RobotPush
-MountainCar
-Figure 2: MuJoCo, robotic and physical simulation tasks with their objectives in our experiments. Detailed descriptions of each task are provided in Appendix G.
+To ensure robust validation, we run $10$ instances per setting with different random seeds and report the average results.
+We also maintain consistent hyperparameters and network architectures across all tasks, detailed in Appendix A.6.
+Figure 2: MuJoCo, robotic, Atari games and physical simulation tasks in our experiments. Detailed descriptions and the environmental reward models of each task are provided in Appendix A.7.
 5.1 Comparison and Discussion
 Baselines.
-We compare SASR with eight baselines to benchmark its performance: (a) RL with an Assistant Reward Agent (ReLara) (Ma et al., 2024b, ), (b) RL Optimizing Shaping Algorithm (ROSA) (Mguni et al.,, 2023), (c) Exploration-Guided RS (ExploRS) (Devidze et al.,, 2022), (d) Count-based static hashing exploration (#Explo) (Tang et al.,, 2017), (e) Random Network Distillation (RND) (Burda et al.,, 2018), (f) Soft Actor-Critic (SAC) (Haarnoja et al., 2018a, ), (g) Twin Delayed DDPG (TD3) (Fujimoto et al.,, 2018), and (h) Proximal Policy Optimization (PPO) (Schulman et al.,, 2017).
-Algorithms (a) to (e) are all reward shaping methods, involving either exploration bonuses or auxiliary agents to shape rewards, while algorithms (f) to (h) are advanced RL algorithms.
-Figure 3 and Table 1 report the average episodic returns and standard errors throughout training.
+We compare SASR with ten baselines to benchmark its performance: (a) the online Distributional Random Network Distillation (DRND) (Yang et al., 2024b), (b) RL with an Assistant Reward Agent (ReLara) (Ma et al., 2024a), (c) General Function Approximation Reward-Free Exploration (GFA-RFE) (Zhang et al., 2024), (d) RL Optimizing Shaping Algorithm (ROSA) (Mguni et al., 2023), (e) Exploration-Guided RS (ExploRS) (Devidze et al., 2022), (f) Count-based static hashing exploration (#Explo) (Tang et al., 2017), (g) Random Network Distillation (RND) (Burda et al., 2018), (h) Soft Actor-Critic (SAC) (Haarnoja et al., 2018a), (i) Twin Delayed DDPG (TD3) (Fujimoto et al., 2018), and (j) Proximal Policy Optimization (PPO) (Schulman et al., 2017).
+Algorithms (a) to (g) are all reward shaping methods, incorporating either exploration bonuses or auxiliary agents to shape rewards, while algorithms (h) to (j) are advanced RL algorithms.
+Figure 3 shows the learning performance of SASR compared with the baselines, while Table 1 reports the average episodic returns with standard errors achieved by the final models over $100$ episodes.
 Our findings indicate that SASR surpasses the baselines in terms of sample efficiency, learning stability, and convergence speed.
-The primary challenge in these environments is the extremely sparse reward given after a long horizon, making exploration essential for timely obtaining successful trajectories.
-Although the exploration strategies of algorithms such as ExploRS, #Explo, and RND are designed to reward novel states, effectively expanding the early exploration space with the direct additional target, they continue focusing on discovering novel states, overlooking the implicit values of these states, which makes them fail to return to the final objectives.
+The primary challenge in these environments is the extremely sparse reward given after a long horizon, making exploration crucial for obtaining successful trajectories in a timely manner.
+Although exploration strategies of algorithms such as ExploRS, #Explo, and RND are designed to reward novel states, effectively expanding the early exploration space with direct additional targets, they continue prioritizing novelty, overlooking the implicit values of these states.
+As a result, they fail to return to the final objectives.
 Figure 3: The learning performance of SASR compared with the baselines.
-Table 1: The average episodic returns and standard errors of SASR and the baselines.
-Algorithms AntStand AntSpeed AntFar AntVeryFar WalkerKeep HumanStand HumanKeep RobotReach RobotPush MountainCar
-SASR 39.12$\pm$2.86 0.94$\pm$0.07 73.92$\pm$4.97 78.64$\pm$3.92 158.24$\pm$5.59 42.63$\pm$2.17 180.98$\pm$4.40 81.29$\pm$6.52 137.06$\pm$12.66 0.91$\pm$0.04
-ReLara 28.66$\pm$1.82 0.33$\pm$0.02 67.77$\pm$4.30 64.07$\pm$4.17 77.14$\pm$8.77 29.72$\pm$1.85 160.31$\pm$7.30 103.56$\pm$7.18 58.71$\pm$6.98 0.89$\pm$0.01
-ROSA 3.80$\pm$0.03 0.02$\pm$0.00 4.71$\pm$0.28 0.64$\pm$0.07 32.14$\pm$1.19 8.55$\pm$0.03 152.38$\pm$4.98 0.27$\pm$0.03 $0$0.90$\pm$0.02
-ExploRS 4.52$\pm$0.04 0.01$\pm$0.00 5.42$\pm$0.22 1.67$\pm$0.10 2.47$\pm$0.13 8.63$\pm$0.03 158.09$\pm$4.42 0.79$\pm$0.04 0.20$\pm$0.08 -0.99$\pm$0.02
-#Explo 6.79$\pm$0.50 0.00$\pm$0.00 12.40$\pm$1.66 0.55$\pm$0.09 131.56$\pm$5.40 28.73$\pm$1.79 160.60$\pm$7.04 4.19$\pm$0.42 6.31$\pm$0.85 0.79$\pm$0.02
-RND 4.23$\pm$0.03 0.02$\pm$0.00 6.66$\pm$0.25 1.00$\pm$0.09 44.78$\pm$1.39 8.67$\pm$0.03 159.79$\pm$4.27 28.18$\pm$2.53 0.04$\pm$0.04 0.94$\pm$0.00
-SAC 15.93$\pm$0.69 0.00$\pm$0.00 9.64$\pm$0.57 20.73$\pm$2.87 70.96$\pm$8.10 9.31$\pm$0.05 4.59$\pm$0.84 45.03$\pm$4.92 0.55$\pm$0.21 -0.05$\pm$0.02
-TD3 0.00$\pm$0.00 0.00$\pm$0.00 0.81$\pm$0.02 0.81$\pm$0.02 18.62$\pm$0.75 5.72$\pm$0.04 0.55$\pm$0.03 0.00$\pm$0.00 0.00$\pm$0.00 0.00$\pm$0.00
-PPO 4.54$\pm$0.04 0.00$\pm$0.00 6.33$\pm$0.24 1.80$\pm$0.11 33.77$\pm$1.11 8.36$\pm$0.03 138.13$\pm$12.64 79.52$\pm$10.80 0.00$\pm$0.00 0.93$\pm$0.00
+Table 1: The average episodic returns and standard errors of all models tested over 100 episodes.
+Tasks SASR DRND-online ReLara GFA-RFE ROSA ExploRS #Explo RND SAC TD3 PPO
+AntStand 94.9$\pm$0.0 67.3$\pm$0.0 90.5$\pm$1.7 54.2$\pm$0.0 3.8$\pm$0.4 5.1$\pm$0.4 17.9$\pm$0.0 4.0$\pm$0.2 31.6$\pm$0.0 0.0$\pm$0.0 4.9$\pm$0.1
+AntFar 139.8$\pm$0.0 93.2$\pm$0.0 115.7$\pm$0.0 86.4$\pm$0.0 1.0$\pm$0.0 12.0$\pm$4.2 75.1$\pm$0.0 4.6$\pm$1.6 25.3$\pm$0.0 1.0$\pm$0.0 7.8$\pm$0.0
+HumanStand 79.8$\pm$2.0 50.6$\pm$0.0 76.2$\pm$0.7 58.2$\pm$0.0 8.8$\pm$0.0 9.3$\pm$0.0 72.7$\pm$0.0 9.3$\pm$0.1 9.9$\pm$0.0 5.5$\pm$0.0 9.0$\pm$0.1
+HumanKeep 195.8$\pm$0.0 154.5$\pm$0.0 194.9$\pm$0.0 141.5$\pm$0.0 169.7$\pm$0.0 182.8$\pm$0.0 195.0$\pm$0.0 180.7$\pm$0.0 2.5$\pm$0.0 1.0$\pm$0.0 138.1$\pm$0.0
+RobotReach 170.2$\pm$0.0 99.8$\pm$0.0 187.9$\pm$0.0 42.1$\pm$0.0 0.1$\pm$0.0 0.7$\pm$0.0 4.6$\pm$0.0 69.3$\pm$0.0 156.5$\pm$0.0 0.0$\pm$0.0 79.5$\pm$0.0
+RobotSlide 132.3$\pm$1.3 127.2$\pm$0.0 111.6$\pm$2.0 115.8$\pm$2.0 11.2$\pm$0.9 4.3$\pm$0.1 3.5$\pm$0.0 4.8$\pm$0.2 0.7$\pm$0.2 0.5$\pm$0.4 0.2$\pm$0.2
+RobotPush 167.1$\pm$0.0 122.2$\pm$0.0 166.9$\pm$0.0 49.1$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 3.7$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0
+RobotPickPlace 1.0$\pm$0.0 1.0$\pm$0.0 1.0$\pm$0.0 0.5$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0
+Pitfall 93.0$\pm$0.0 92.0$\pm$0.0 40.3$\pm$0.0 89.4$\pm$0.0 0.0$\pm$0.0 57.6$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 4.6$\pm$0.0 0.5$\pm$0.0 0.0$\pm$0.0
+Frogger 14.2$\pm$0.0 11.7$\pm$0.0 11.6$\pm$0.0 7.9$\pm$0.0 9.8$\pm$0.0 8.3$\pm$0.0 11.9$\pm$0.0 10.5$\pm$0.0 0.8$\pm$0.0 0.7$\pm$0.0 0.0$\pm$0.0
+Montezuma 6737.9$\pm$0.0 6828.5$\pm$0.0 2421.9$\pm$0.0 4755.3$\pm$0.0 4294.4$\pm$0.0 3971.5$\pm$0.0 1400.1$\pm$0.0 5494.3$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0 0.0$\pm$0.0
+Solaris 42.1$\pm$0.0 21.3$\pm$0.7 20.3$\pm$0.0 26.3$\pm$0.0 0.1$\pm$0.0 17.0$\pm$0.0 1.2$\pm$0.8 9.8$\pm$0.0 6.0$\pm$0.0 0.4$\pm$0.0 1.5$\pm$0.0
+Freeway 22.4$\pm$0.0 19.8$\pm$0.0 21.5$\pm$0.0 10.1$\pm$0.0 18.0$\pm$0.0 17.5$\pm$0.0 6.9$\pm$0.0 13.0$\pm$0.0 0.1$\pm$0.0 0.2$\pm$0.0 0.0$\pm$0.0
+MountainCar 1.0$\pm$0.0 1.0$\pm$0.0 1.0$\pm$0.0 1.0$\pm$0.0 -0.9$\pm$0.0 -1.0$\pm$0.0 1.0$\pm$0.0 1.0$\pm$0.0 -0.1$\pm$0.0 0.0$\pm$0.0 0.9$\pm$0.0
 SASR outperforms the baselines primarily due to its self-adaptive reward evolution mechanism.
-In the early phases, SASR encourages exploration by injecting substantial random rewards to optimize the agent in multiple directions, thereby increasing the likelihood of collecting positive samples.
-Unlike novelty-driven exploration strategies that reward only novel states, SASR relaxes this assumption by rewarding random states, under the belief that all states are worth exploring, and this approach has proven empirically effective.
-Additionally, SASR leverages the success rate as auxiliary information to assess each state’s contribution to task completion, aligning with the agent’s primary objectives.
-As more data is collected, the success rate becomes more accurate, and the shaped reward provides more valuable guidance, significantly enhancing the exploitation and stabilizing the convergence toward optimal policies.
-Cooperatively, these strategies improve SASR’s sample efficiency and convergence stability in challenging tasks.
-While ReLara used a similar exploration mechanism to SASR by perturbing reward functions, it relies on an independent black-box agent, requiring more iterations to converge.
-In contrast, SASR’s success rate sampling is more direct, reducing the delay in converging to valuable information.
+In the early phases, SASR encourages exploration by injecting substantial random rewards, optimizing the agent in multiple directions, and increasing the likelihood of collecting positive samples.
+Moreover, since most states are initially classified as failures, their success rates decrease.
+As a result, unvisited states receive relatively higher rewards, encouraging further exploration.
+This mechanism resembles intrinsic motivation, assigning higher rewards to novel states, and effectively guiding the agent to expand the exploration space.
+As more data is collected, the success rate estimation becomes more accurate, the shaped reward provides more reliable guidance, enhancing exploitation and stabilizing convergence.
+Together, these strategies improve SASR’s sample efficiency and convergence stability in challenging tasks.
+While ReLara used a similar exploration mechanism by perturbing reward functions, it relies on an independent black-box agent, requiring more iterations to converge.
+In contrast, SASR’s success rate sampling is more direct, reducing delays in acquiring valuable information.
 ReLara’s advantage lies in incorporating the policy agent’s actions into reward construction, as seen in the RobotReach task, where the target point is randomly selected in each episode.
 In this case, ReLara outperforms SASR due to access to action information.
-This can also be achieved in SASR by treating actions as additional features in the state vector.
+However, SASR can achieve the same by incorporating actions as additional features in the state vector.
 5.2 Effect of Self-Adaptive Success Rate Sampling
-SASR introduces a novel self-adaptive mechanism that balances exploration and exploitation by maintaining the randomness of the shaped rewards.
-To further explore the effect of this mechanism, we use the AntStand task as a case study, analyzing the shaped rewards learned at different stages of training.
-Figure 4 (bottom) shows the learning curve, while Figure 4 (top) illustrates the distributions of generated rewards over the “height of the ant" feature, one dimension in the state vector.
+SASR introduces a novel self-adaptive mechanism that balances exploration and exploitation by modulating the randomness of the shaped rewards.
+To further investigate the effect of this mechanism, we use the AntStand task as a case study, analyzing the shaped rewards learned at different training stages.
+Figure 4 (bottom) shows the learning curve, while Figure 4 (top) illustrates the distributions of generated rewards over the “height of the ant” feature, a dimension in the state vector.
 Figure 4: Distributions of the shaped rewards over the height of the ant robot in the AntStand task at different training stages.
@@
 Figure 5: The density of visited states in the MountainCar task for four training periods.
-As learning progresses, the shaped rewards demonstrate two key attributes: the values transition from random to meaningful, and the variance decreases from uncertain to deterministic.
-Besides, although the sampled rewards fluctuate, their center gradually shows a positive linear correlation with the robot’s height.
-In early phases, the shaped rewards contain significant randomness due to higher uncertainty.
-Although these random signals offer limited information, they drive the agent to take tiny optimization steps in diverse directions, effectively shifting the anchors of the policies, thus broadens the actions sampled from SAC’s stochastic policy, encouraging exploration and increasing the diversity of collected samples.
-In later phases, the rewards converge to closely match the height of the robot, a meaningful and intuition-aligned metric, which enhances the agent’s exploitation.
-To further investigate the exploration behavior of SASR, we compare the density of the visited states throughout the training in the MountainCar task with four representative exploration strategies: (1) ReLara, perturbating on both rewards and actions; (2) #Explo, rewarding novel states; (3) SAC, entropy-regularized exploration; and (4) NoisyNet (Fortunato et al.,, 2018), perturbing network weights.
-The density for every 25k steps is shown in Figure 5.
+As learning progresses, the shaped rewards exhibit two key attributes: values transition from random to meaningful, and variance decreases from uncertain to deterministic.
+Although the sampled rewards fluctuate, their means gradually show a positive linear correlation with the robot’s height.
+In the early phases, the shaped rewards contain significant randomness due to high uncertainty.
+While these random signals offer limited information, they drive the agent to take small optimization steps in diverse directions, effectively shifting the policy anchors, expanding the action space sampled from SAC’s stochastic policy, promoting exploration, and increasing sample diversity.
+In later phases, rewards stabilizes, closely aligning with the robot’s height – a meaningful and intuitive metric – enhancing the agent’s exploitation.
+To further investigate SASR’s exploration behavior, we compare the visited state density throughout training in the MountainCar task with five representative exploration strategies: (1) ReLara, which perturbs both rewards and actions; (2) #Explo and (3) RND, which rewards novel states; (4) SAC, which uses entropy-regularized exploration; and (5) NoisyNet (Fortunato et al., 2018), which perturbs network weights.
+The state density for every 25k steps is shown in Figure 5.
 We observe that SASR progressively covers a wider range of the state space.
-From 50k to 100k steps, SASR reaches positions near the goal, driven by the success rate mechanism.
-In contrast, ReLara covers a similar range to SASR, but is less smooth and takes longer to reach the right side.
-#Explo shows no clear rightward shift, as it focuses on rewarding novelty without considering the inherent value of states.
-SAC’s exploration is relatively narrow, making it easier to trap in local optima.
-NoisyNet shows a narrowing range as network weight perturbations diminish through optimization.
+From 50k to 100k steps, SASR reaches positions near the goal, driven by its success rate mechanism.
+In contrast, ReLara and RND cover similar ranges to SASR, but are less smooth and take longer to reach the right side.
+#Explo shows no clear rightward shift, as it prioritizes novelty and ignores the inherent value of states.
+SAC’s exploration is relatively narrow, making it prone to getting trapped in local optima.
+NoisyNet’s range narrows over time as perturbations diminish through optimization.
 Overall, SASR demonstrates more effective exploration and collects valuable samples sooner, leading to faster convergence.
 5.3 Ablation Study
-We conduct ablation studies to investigate key components of SASR. We select six representative tasks to report the experimental results, and the quantitative data is provided in Appendix D.
+We conduct ablation studies to investigate key components of SASR. We select six representative tasks and report the experimental results, with quantitative data provided in Appendix A.5.
 (a) The impact of omitting Beta distribution sampling on the performance of SASR.
-(b) Learning performance of SASR with different retention rates $\phi$.
-(c) Learning performance of SASR with different bandwidths $h$ of Gaussian kernel.
+(b) Learning performance of SASR with shaped reward function $R^{S}(s,a)$ and $R^{S}(s)$.
+(c) Learning performance of SASR with different retention rates $\phi$.
 (d) Learning performance of SASR with different RFF feature dimensions $M$.
@@
 (Figure 6(a)) We examine a variant of SASR that omits Beta distribution sampling, instead directly using the success rate $N_{S}(s_{i})/(N_{S}(s_{i})+N_{F}(s_{i}))$.
-In the early stages, limited experience makes this success rate an unreliable estimate, and using this fixed, overly confident value can mislead the agent.
-Furthermore, skipping Beta distribution sampling removes the exploration driven by random rewards, leading to narrower exploration.
-The results highlight the critical role of Beta distribution sampling for effective learning.
+In the early stages, limited experience makes this success rate an unreliable estimate, and using a fixed, overly confident value can mislead the agent.
+Furthermore, skipping Beta distribution sampling eliminates exploration driven by random rewards, leading to narrower exploration.
+The results highlight the critical role of Beta distribution sampling in effective learning.
+Reward function over state-action pair.
+(Figure 6(b)) We extend SASR with a reward function over state-action pairs, $r(s,a)$.
+The comparison results show that both settings perform similarly.
+However, encoding actions into the reward function increases dimensionality, complicating density estimation and correlation assessment.
+Furthermore, the state and action vectors may have different distributions, potentially reducing KDE estimation accuracy.
 Retention rate $\phi$.
-(Figure 6(b)) The retention rate controls the estimation of counts, which affects the confidence of Beta distributions.
-A high retention rate ($\phi=1$) preserves all samples, resulting in a densely populated and redundant state pool, causing the Beta distribution to be prematurely overconfident, which degrades performance.
-Conversely, a low retention rate ($\phi=0.01$) slows convergence by requiring more iterations to gather sufficient samples.
-The results suggest that an appropriate retention rate is important.
-Bandwidth $h$ of Gaussian kernel.
-(Figure 6(c)) The bandwidth $h$ controls the smoothness of the kernel functions.
-Beyond fixed bandwidths, we also test a linearly decreasing configuration ($h:0.5\rightarrow 0.1$), which reflects increasing confidence in KDE.
-Results indicate that a small bandwidth ($h=0.01$) increases the distance between samples, causing many to have zero estimated density, while a large bandwidth ($h=1$) makes samples indistinguishable due to an overly flat kernel function.
-Both cases result in suboptimal performance.
-The decreasing bandwidth setting offers no significant improvement and tends to reduce stability due to inconsistent density estimations.
+(Figure 6(c)) The retention rate directly influences the confidence level of Beta distributions.
+A high retention rate ($\phi=1$) preserves all samples, resulting in a densely populated, redundant state pool, which makes the Beta distribution prematurely overconfident and degrades performance.
+Conversely, a low retention rate ($\phi=0.01$) slows convergence as more iterations are required to gather sufficient samples.
+The results suggest that an appropriate retention rate is crucial.
 RFF feature dimensions $M$.
-(Figure 6(d)) SASR shows relatively low sensitivity to $M$, provided it is large enough to capture the states’ complexity.
-Results show that values like $M=500,1000,2000$ all yield similar performance levels, while significantly low dimensions, such as $M=50$, degrade performance.
-This is also suggested in the original RFF study (Rahimi and Recht,, 2007).
+(Figure 6(d)) SASR shows relatively low sensitivity to $M$, provided it is sufficiently large to capture the complexity of the states.
+Results show that values like $M=500,1000,2000$ all yield similar performance, while significantly lower dimensions, such as $M=50$, degrade performance.
+These findings align with the original RFF study (Rahimi & Recht, 2007).
 Shaped reward weight factor $\lambda$.
-(Figure 6(e)) Regarding the shaped reward weight factor $\lambda$, we observe that SASR performs better with intermediate values like $\lambda=0.4,0.6,0.8$.
+(Figure 6(e)) SASR performs better with intermediate values like $\lambda=0.4,0.6,0.8$.
 At $\lambda=0.2$, the minimal shaped reward scale reduces state differentiation, leading to suboptimal performance.
-At $\lambda=1$, aligning the shaped reward scale with the environmental reward introduces excessive uncertainty and randomness, which may overwhelm the feedback and hinder learning.
-The findings highlight that maintaining a balanced reward scale is important for optimal learning outcomes.
+At $\lambda=1$, aligning the shaped reward scale with the environmental reward introduces excessive uncertainty and randomness, potentially overwhelming feedback and hindering learning.
+The findings emphasize that maintaining a balanced reward scale is important for optimal learning outcomes.
 6 Conclusion and Discussion
-In this paper, we propose SASR, a self-adaptive reward shaping algorithm based on success rates, to address the sparse-reward challenge.
-SASR achieves an exploration and exploitation balance mechanism by generating shaped rewards from evolving Beta distributions.
-Experiments demonstrate that this adaptability significantly enhances the agent’s convergence speed.
-Additionally, the implementation of KDE and RFF formulates a highly efficient, non-parametric, and learning-free approach to deriving Beta distributions.
-This mechanism also provides a sound alternative to traditional count-based RS strategies, adapting effectively to continuous environments.
+In this paper, we propose SASR, a self-adaptive reward shaping algorithm based on success rates to address the sparse-reward challenge.
+SASR achieves a balance between exploration and exploitation by generating shaped rewards from evolving Beta distributions.
+Experiments demonstrate that this adaptability significantly improves the agent’s convergence speed.
+Additionally, the use of KDE and RFF provides a computationally efficient approach to deriving Beta distributions.
+This mechanism also offers a sound alternative to traditional count-based RS strategies, adapting effectively to continuous environments.
 Our evaluations confirm the superior performance of SASR in terms of sample efficiency and learning stability.
-While SASR is designed for sparse-reward environments, in settings with dense rewards, the introduction and maintenance of additional shaped rewards could be unnecessary.
-Exploring the extension of SASR to dense reward scenarios presents a promising direction for further research.
-Moreover, the derivation of Beta distributions depends heavily on the samples stored in the success and failure buffers.
-Currently, our method does not account for the relationships or varying importance of different states within the same trajectory, making the algorithm sensitive to the retention rate $\phi$.
-Therefore, developing an adaptive retention rate or better mechanisms for managing the buffers are crucial avenues for future improvement.
-7 Acknowledgement
-Appendix A Derivation of Random Fourier Features
-We incorporate Random Fourier Features (RFF) (Rahimi and Recht,, 2007) to approximate the kernel functions in the KDE process for the SASR algorithm.
+While SASR is designed for sparse-reward environments, in dense-reward settings, the additional shaped rewards may be unnecessary.
+Extending SASR to such scenarios presents a promising direction for further research.
+Moreover, the derivation of Beta distributions relies on the samples stored in the success and failure buffers.
+Currently, our method does not consider the relationships or varying importance of different states within the same trajectory, making it sensitive to the retention rate.
+Therefore, developing an adaptive retention rate or improved buffer management mechanisms is crucial for future improvement.
+Acknowledgement
+This research is supported by the National Research Foundation Singapore and DSO National Laboratories under the AI Singapore Programme (Award Number: AISG2-RP-2020-016).
+This research is also partially supported by an Academic Research Grant No.
+MOE-T2EP20121-0015 from the Ministry of Education in Singapore.
+Appendix A Appendix
+A.1 Derivation of Random Fourier Features
+We incorporate Random Fourier Features (RFF) (Rahimi & Recht, 2007) to approximate the kernel functions in the KDE process for the SASR algorithm.
 Let the original state be $k$-dimensional, denoted as $\bm{s}\in\mathbb{R}^{k}$, and the kernel function be $k(\bm{s}_{i},\bm{s}_{j})$.
@@
 The RFF-based kernel function can be approximated by the inner product of the mapped features in the $M$-dimensional space.
-Appendix B Derivation of Computation Complexity
+A.2 Computational Complexity
 In this section, we derive the computational complexity to retrieve the success or failure counts $N_{S}$ and $N_{F}$ for each iteration.
@@
 For each state, the mapping function calculation involves:
+  1. 1.
     Matrix multiplication $\bm{W}^{T}\bm{s}$: $kM$.
+  2. 2.
     Addition $\bm{W}^{T}\bm{s}+\bm{b}$: $M$.
+  3. 3.
     Cosine calculation $\cos(\bm{W}^{T}\bm{s}+\bm{b})$: $M$.
@@
 Since the first two terms $O(DkM)$ and $O(BkM)$ are dominated by the last term $O(MDB)$ when the buffer size and the batch size are large, the overall computational complexity to retrieve the corresponding counts can be approximated as $O(MDB)$.
-Appendix C Comparison of SASR with and without RFF
+A.3 Experiments on Time and Space Complexity
+A.3.1 Time and Space Complexity Comparison
+In this section, we analyze the time and space overhead introduced by SASR and other representative reward-shaping methods.
+Below, we summarize the computational and memory costs of the RS baselines, introduced by the shaped reward generation.
+  • SASR (ours) calculates shaped rewards using RFF, which essentially is matrix operations, without additional networks/models learning processes.
+Regarding the memory costs, the buffers $D_{S}$ and $D_{F}$ are much smaller than the replay buffer used in the backbone SAC algorithm, due to the retention rate $\phi$.
+While considering the scalability for larger problems, we have implemented an alternative approach by augmenting the original replay buffer in the backbone SAC algorithm with a success or failure flag.
+This approach avoids the need for additional buffers.
+  • ReLara (Ma et al., 2024a) requires an additional RL agent (of the same scale as the original RL agent) and an additional replay buffer.
+  • ROSA (Mguni et al., 2023) involves a competition agent (the same sacle as the original RL agent) and a switching model (a neural network).
+  • ExploRS (Devidze et al., 2022) requires learning two parameterized networks: one for a self-supervised reward model and another for the exploration bonus.
+  • #Explo (Tang et al., 2017) requires a hash function to discretize the state space and a hash table to store the state-visitation counts.
+  • RND (Burda et al., 2018) uses a random network distillation module to compute the intrinsic rewards.
+Furthermore, we report the computational and memory costs of SASR and the RS baselines in two tasks: AntStand and Frogger, the results are shown in Table 3 and Table 4, respectively.
+To provide a more intuitive comparison, we report the relative value normalized to our SASR, in this case, if the value $>1$, it indicates that the baseline is more computationally or memory expensive than SASR, and vice versa.
+Table 3: Average maximum memory consumption during the training process, normalized to SASR.
+Tasks SASR ReLara ROSA ExploRS #Explo RND
+AntStand 1 3.67 4.12 2.05 0.89 0.12
+Frogger 1 5.21 4.33 2.64 0.92 0.09
+Table 4: Average training time, normalized to SASR.
+Tasks SASR ReLara ROSA ExploRS #Explo RND
+AntStand 1 1.87 2.12 1.67 1.08 1.11
+Frogger 1 1.98 3.17 1.72 1.24 1.06
+A.3.2 Comparison of SASR with and without RFF
 To evaluate the effect of introducing RFF, we compare the training time of SASR with and without RFF, also with the backbone SAC algorithm, the results are shown in Table 5.
@@
 SASR w/o RFF 8.12 8.72 8.37 6.53 11.12 9.21
-Appendix D Supplementary Experimental Results for Ablation Study
-We provide the detailed quantitative results of the ablation study in this section.
+A.4 Auto Hyperparameter Selection
+To improve the robustness and generalization of the SASR algorithm, we propose some potential autonomous hyperparameter selection strategies, mainly designed for the bandwidth $h$ of the kernel function and the RFF feature dimension $M$.
+For bandwidth $h$, we can use the empirical formula Silverman’s Rule of Thumb (Wilcox, 2011):
+$h=1.06\cdot\sigma\cdot N^{-1/5},$ (26)
+or cross-validation to determine the optimal bandwidth.
+For the RFF dimension $M$, it is directly related to the bandwidth $h$. After determining $h$, we can use the formula mentioned in the RFF theory to determine $M$:
+$M=O(\frac{1}{\epsilon^{2}}\log{\frac{N}{\delta}}),$ (27)
+where $\epsilon$ and $\delta$ are the error and confidence parameters.
+Another method is to compare the Frobenius norm error between the RFF approximated kernel matrix $K^{RFF}$ and the true kernel matrix $K^{Gaussian}$ to select $M$: $\|K^{RFF}-K^{Gaussian}\|_{F}$.
+A.5 Supplementary Experimental Results for Ablation Study
+In this section, we provide the detailed quantitative results of the ablation study.
 Table 6: Ablation study #1: The average episodic returns and standard errors of SASR and the variant without sampling from Beta distributions.
 Tasks SASR (with sampling) SASR (without sampling)
-AntStand $\bm{39.12\pm 2.86}$ $23.25\pm 2.12$
-AntFar $\bm{73.92\pm 4.97}$ $37.86\pm 5.16$
-HumanStand $\bm{42.63\pm 2.17}$ $9.30\pm 0.083$
-HumanKeep $\bm{180.98\pm 4.40}$ $170.51\pm 10.68$
-RobotReach $\bm{81.29\pm 6.52}$ $33.76\pm 4.41$
-RobotPush $\bm{137.06\pm 12.66}$ $70.37\pm 11.97$
-Table 5: Ablation study #2: The average episodic returns and standard errors of SASR with different retention rates.
+AntStand 94.92$\pm$0.00 54.48 $\pm$ 1.29
+AntFar 139.84$\pm$0.00 92.77 $\pm$ 1.53
+HumanStand 79.83$\pm$2.03 9.77 $\pm$ 0.02
+HumanKeep 195.77$\pm$0.00 185.00 $\pm$ 0.00
+RobotReach 170.18$\pm$0.00 110.29 $\pm$ 2.93
+RobotPush 167.14$\pm$0.00 86.82 $\pm$ 0.00
+Table 7: Ablation study #2: The average episodic returns and standard errors of SASR with reward function on state-action pair or state only.
+Tasks SASR (with $R^{S}(s)$) SASR (with $R^{S}(s,a)$)
+AntStand 94.92$\pm$0.00 85.61$\pm$1.30
+AntFar 139.84$\pm$0.00 132.49$\pm$2.92
+HumanStand 79.83$\pm$2.03 78.93$\pm$0.65
+HumanKeep 195.77$\pm$0.00 192.54$\pm$0.16
+RobotReach 170.18$\pm$0.00 151.95$\pm$5.74
+RobotPush 167.14$\pm$0.00 179.76$\pm$1.66
+Table 8: Ablation study #3: The average episodic returns and standard errors of SASR with different retention rates.
 Tasks $\phi=1$ $\phi=0.1$ (default) $\phi=0.01$
-AntStand $17.98\pm 1.86$ $\bm{39.12\pm 2.86}$ $20.27\pm 2.08$
-AntFar $19.95\pm 3.00$ $\bm{73.92\pm 4.97}$ $48.47\pm 5.28$
-HumanStand $9.25\pm 0.08$ $\bm{42.63\pm 2.17}$ $19.17\pm 1.52$
-HumanKeep $179.10\pm 7.06$ $\bm{180.98\pm 4.40}$ $180.13\pm 6.27$
-RobotReach $54.07\pm 8.39$ $\bm{81.29\pm 6.52}$ $35.18\pm 5.95$
-RobotPush $3.91\pm 1.22$ $\bm{137.06\pm 12.66}$ $1.88\pm 0.47$
-Table 6: Ablation study #3: The average episodic returns and standard errors of SASR with different bandwidths $h$ of Gaussian kernel.
-Tasks $h=0.01$ $h=0.1$ $h=0.2$ (default) $h=1$ $h=0.5\rightarrow 0.1$
-AntStand $7.94\pm 0.28$ $16.77\pm 1.18$ $\bm{39.12\pm 2.86}$ $12.76\pm 0.40$ $19.00\pm 1.63$
-AntFar $11.68\pm 0.96$ $44.31\pm 3.34$ $\bm{73.92\pm 4.97}$ $27.96\pm 2.46$ $72.71\pm 4.57$
-HumanStand $9.31\pm 0.06$ $\bm{43.64\pm 1.94}$ $42.63\pm 2.17$ $9.32\pm 0.06$ $13.81\pm 0.93$
-HumanKeep $178.51\pm 4.99$ $\bm{181.44\pm 4.73}$ $180.98\pm 4.40$ $178.22\pm 4.97$ $178.04\pm 5.12$
-RobotReach $37.03\pm 4.19$ $72.30\pm 5.67$ $\bm{81.29\pm 6.52}$ $35.87\pm 4.57$ $49.47\pm 5.1$
-RobotPush $0.40\pm 0.13$ $8.80\pm 4.44$ $\bm{137.06\pm 12.66}$ $0.12\pm 0.06$ $0.34\pm 0.13$
+AntStand 45.71$\pm$7.57 94.92$\pm$0.00 62.85$\pm$3.49
+AntFar 70.07$\pm$3.30 139.84$\pm$0.00 103.65$\pm$2.57
+HumanStand 9.88$\pm$0.01 79.83$\pm$2.03 66.46$\pm$2.96
+HumanKeep 195.00$\pm$0.00 195.77$\pm$0.00 194.77$\pm$0.10
+RobotReach 154.32$\pm$0.89 170.18$\pm$0.00 112.46$\pm$0.90
+RobotPush 2.96$\pm$1.97 167.14$\pm$0.00 1.75$\pm$1.24
 Table 9: Ablation study #4: The average episodic returns and standard errors of SASR with different RFF feature dimensions $M$.
 Tasks $M=50$ $M=500$ $M=1000$ (default) $M=2000$
-AntStand $3.56\pm 0.12$ $12.08\pm 1.36$ $39.12\pm 2.86$ $\bm{42.52\pm 3.14}$
-AntFar $27.76\pm 3.24$ $36.31\pm 2.69$ $\bm{73.92\pm 4.97}$ $61.79\pm 4.63$
-HumanStand $9.46\pm 0.05$ $25.42\pm 1.83$ $42.63\pm 2.17$ $\bm{52.70\pm 2.13}$
-HumanKeep $178.51\pm 4.84$ $181.81\pm 4.60$ $\bm{180.98\pm 4.40}$ $180.58\pm 4.63$
-RobotReach $65.39\pm 6.22$ $38.25\pm 4.37$ $\bm{81.29\pm 6.52}$ $45.98\pm 4.90$
-RobotPush $33.58\pm 4.86$ $54.56\pm 6.66$ $\bm{137.06\pm 12.66}$ $45.15\pm 6.15$
+AntStand 5.21$\pm$0.45 50.68$\pm$6.40 94.92$\pm$0.00 96.80$\pm$8.42
+AntFar 98.88$\pm$2.64 72.17$\pm$5.07 139.84$\pm$0.00 129.87$\pm$0.63
+HumanStand 9.87$\pm$0.01 78.82$\pm$0.52 79.83$\pm$2.03 77.73$\pm$1.47
+HumanKeep 193.84$\pm$0.61 194.71$\pm$0.15 195.77$\pm$0.00 195.86$\pm$0.03
+RobotReach 119.09$\pm$26.92 122.89$\pm$4.51 170.18$\pm$0.00 94.87$\pm$15.56
+RobotPush 71.67$\pm$27.53 161.70$\pm$11.54 167.14$\pm$0.00 150.20$\pm$8.70
 Table 10: Ablation study #5: The average episodic returns and standard errors of SASR with different shaped reward weight factors.
 Tasks $\lambda=0.2$ $\lambda=0.4$ $\lambda=0.6$ (default) $\lambda=0.8$ $\lambda=1.0$
-AntStand $7.60\pm 0.73$ $17.25\pm 1.58$ $\bm{39.12\pm 2.86}$ $28.43\pm 2.07$ $2.82\pm 0.09$
-AntFar $22.51\pm 2.60$ $34.04\pm 3.87$ $\bm{73.92\pm 4.97}$ $67.62\pm 4.89$ $18.05\pm 1.79$
-HumanStand $9.53\pm 0.04$ $33.95\pm 1.90$ $\bm{42.63\pm 2.17}$ $30.65\pm 1.65$ $12.23\pm 0.42$
-HumanKeep $\bm{182.57\pm 4.32}$ $180.22\pm 4.60$ $180.98\pm 4.40$ $179.74\pm 4.53$ $179.62\pm 4.70$
-RobotReach $41.05\pm 4.17$ $\bm{82.33\pm 6.37}$ $81.29\pm 6.52$ $71.30\pm 6.54$ $15.89\pm 2.06$
-RobotPush $9.18\pm 1.99$ $118.92\pm 10.79$ $\bm{137.06\pm 12.66}$ $65.51\pm 12.71$ $47.98\pm 9.64$
-Appendix E Network Structures and Hyperparameters
-E.1 Network Structures
+AntStand 35.71$\pm$0.92 59.82$\pm$2.71 94.92$\pm$0.00 75.61$\pm$1.41 3.16$\pm$0.35
+AntFar 99.83$\pm$3.25 119.82$\pm$1.18 139.84$\pm$0.00 119.35$\pm$1.80 80.71$\pm$4.74
+HumanStand 9.81$\pm$0.02 75.35$\pm$1.06 79.83$\pm$2.03 70.96$\pm$0.51 28.94$\pm$0.46
+HumanKeep 194.68$\pm$0.08 194.21$\pm$0.38 195.77$\pm$0.00 193.85$\pm$0.42 194.89$\pm$0.10
+RobotReach 131.61$\pm$4.51 154.16$\pm$5.20 170.18$\pm$0.00 169.00$\pm$2.56 74.23$\pm$4.23
+RobotPush 13.07$\pm$1.65 193.56$\pm$3.85 167.14$\pm$0.00 178.90$\pm$0.00 192.07$\pm$0.87
+Bandwidth $h$ of Gaussian kernel.
+The bandwidth $h$ controls the smoothness of the kernel functions.
+Beyond fixed bandwidths, we also test a linearly decreasing configuration ($h:0.5\rightarrow 0.1$), which reflects increasing confidence in KDE.
+Results indicate that a small bandwidth ($h=0.01$) increases the distance between samples, causing many to have zero estimated density, while a large bandwidth ($h=1$) makes samples indistinguishable due to an overly flat kernel function.
+Both cases result in suboptimal performance.
+The decreasing bandwidth setting offers no significant improvement and tends to reduce stability due to inconsistent density estimations.
+Table 11: Ablation study #6: The average episodic returns and standard errors of SASR with different bandwidths $h$ of Gaussian kernel.
+Tasks $h=0.01$ $h=0.1$ $h=0.2$ (default) $h=1$ $h=0.5\rightarrow 0.1$
+AntStand 10.71$\pm$2.52 57.22$\pm$3.87 94.92$\pm$0.00 17.74$\pm$2.53 68.40$\pm$1.54
+AntFar 17.58$\pm$2.84 99.80$\pm$4.41 139.84$\pm$0.00 25.83$\pm$8.62 136.49$\pm$4.15
+HumanStand 9.89$\pm$0.01 64.47$\pm$1.87 79.83$\pm$2.03 9.90$\pm$0.02 58.79$\pm$2.76
+HumanKeep 194.92$\pm$0.02 194.00$\pm$0.57 195.77$\pm$0.00 193.06$\pm$0.46 194.59$\pm$0.18
+RobotReach 128.57$\pm$3.83 97.35$\pm$19.12 170.18$\pm$0.00 134.02$\pm$2.02 59.39$\pm$26.02
+RobotPush 2.29$\pm$1.62 122.45$\pm$37.58 167.14$\pm$0.00 0.00$\pm$0.00 0.01$\pm$0.01
+A.6 Network Structures and Hyperparameters
+A.6.1 Network Structures
 Figure 7 illustrates the structures of the policy network and Q-network employed in our experiments.
@@
 Figure 7: The structures of policy network and Q-network in our implementation.
-E.2 Hyperparameters
+A.6.2 Hyperparameters
 We have observed that SASR demonstrated high robustness and was not sensitive to hyperparameter choices. Table 12 shows the set of hyperparameters that we used in all of our experiments.
@@
 burn-in steps $5000$
-Appendix F Compute Resources
+A.6.3 Compute Resources
 The experiments in this paper were conducted on a computing cluster, with the detailed hardware configurations listed in Table 13.
@@
 Brand Supermicro 2022
-Appendix G Configurations of Tasks
-We provide the detailed configurations of the tasks in the experiments in this section.
+A.7 Configurations of Tasks
+In this section, we provide the detailed configurations of the tasks in the experiments.
   • AntStand: The ant robot is trained to stand over a target position. The reward is given if the ant robot reaches the target height. Maximum episode length is 1000 steps.
-  • AntSpeed: The ant robot is trained to move as fast as possible. The reward is given if the ant robot reaches a target speed. Maximum episode length is 1000 steps.
   • AntFar: The ant robot is trained to reach a target position far from the starting point. The reward is given if the ant robot reaches the target position. Maximum episode length is 1000 steps.
-  • AntVeryFar: The ant robot is trained to reach a target position very far from the starting point.
-The reward is given if the ant robot reaches the target position.
-Maximum episode length is 1000 steps.
-  • WalkerKeep: The walker robot is trained to keep a target height.
-The robot is initialized by standing, and the reward is given if the robot maintains the target height.
-Maximum episode length is 1000 steps.
   • HumanStand: The human robot is trained to stand over a target position.
@@
 Maximum episode length is 500 steps.
+  • RobotSlide: The robot arm is trained to slide an object to a target position.
+The target position is randomly generated on the table, and the reward is given if the object reaches the target position.
+Maximum episode length is 500 steps.
+  • RobotPickPlace: The robot arm is trained to pick and place an object to a target position.
+The target position is randomly generated in the space, and the reward is given if the object reaches the target position.
+Maximum episode length is 500 steps.
+  • Pitfall: The agent is tasked with collecting all the treasures in a jungle while avoiding the pitfalls.
+The reward is given if the agent collects one treasure, while if the agent falls into a pitfall, the episode ends.
+Maximum episode length is 2000 steps.
+  • Frogger: The agent is trained to cross frogs on a river.
+The reward is given when each frog is crossed, and the episode ends if all frogs are crossed or fall into the river.
+Maximum episode length is 2000 steps.
+  • MontezumaRevenge: The agent is trained to navigate through a series of rooms to collect keys and reach the final room.
+The reward is given if the agent successfully reaches one new room.
+Maximum episode length is 5000 steps.
+  • Solaris: The agent controls a spaceship to blast enemies and explore new galaxies.
+The reward is given if the agent destroys one enemy spaceship and enters a new galaxy.
+Maximum episode length is 2000 steps.
+  • Freeway: The agent is trained to guide the chicken across multiple lanes of heavy traffic.
+The reward is given if one chicken crosses one lane, while the episode ends if all chickens are crossed or hit by a car.
+Maximum episode length is 2000 steps.
   • MountainCar: The car is trained to reach the top of the right hill. The reward is given if the car reaches the top. Maximum episode length is 1000 steps.
+Furthermore, we provide the detailed dimensions of the states in our evaluated tasks in Table 14.
+Table 14: The dimensions of the states in the evaluated tasks.
+Domain (Tasks) Dimension
+Ant robot (AntStand, AntFar) 105
+Humanoid robot (HumanStand, HumanKeep) 348
+RobotReach 20
+RobotPush, RobotSlide and RobotPickPlace 35
+Atari games (MontezumaRevenge, PitFall, Frooger, Solaris, Freeway) $84\times 84=7056$
+MountainCar 2